### v1

# A Review of Large Language Models and Autonomous Agents in Chemistry

Abstract
Large language models (LLMs) are emerging as a powerful tool in chemistry across multiple domains. In chemistry, LLMs are able to accurately predict properties, design new molecules, optimize synthesis pathways, and accelerate drug and material discovery. A core emerging idea is combining LLMs with chemistry-specific tools like synthesis planners and databases, leading to so-called “agents.” This review covers LLMs’ recent history, current capabilities, design, challenges specific to chemistry, and future directions. Particular attention is given to agents and their emergence as a cross-chemistry paradigm. Agents have proven effective in diverse domains of chemistry, but challenges remain. It is unclear if creating domain-specific versus generalist agents and developing autonomous pipelines versus "co-pilot" systems will accelerate chemistry. An emerging direction is the development of multi-agent systems using a human-in-the-loop approach. Due to the incredibly fast development of this field, a repository has been built to keep track of the latest studies: https://github.com/ur-whitelab/LLMs-in-science.
Keywords
1 Introduction
The integration of Machine Learning (ML) and Artificial Intelligence (AI) into chemistry has spanned several decades1, 2, 3, 4, 5, 6, 7, 8, 9, 10. Although applications of computational methods in quantum chemistry and molecular modeling from the 1950s-1970s were not considered AI, they laid the groundwork. Subsequently in the 1980s expert systems like DENDRAL11, 12 were expanded to infer molecular structures from mass spectrometry data.13 At the same time, Quantitative Structure-Activity Relationship (QSAR) Models were developed5 that would use statistical methods to predict the effects of chemical structure on activity.14, 15, 16, 17 In the 1990s, neural networks were introduced to domains such as drug design, as summarized well by Yang et al. 5 and Goldman and Walters 18, although they were limited by the computational resources of the time. With an explosion of data from High-Throughput Screening (HTS)19, 20, models then started to benefit from vast datasets of molecular structures and their biological activities. Furthermore, ML algorithms such as Support Vector Machines and Random Forests became popular for classification and regression tasks in cheminformatics,1 offering improved performance over traditional statistical methods.21
Deep learning transformed the landscape of ML in chemistry and materials science in the 2010s.22 Recurrent Neural Networks (RNNs),23, 24, 25, 26, 27 Convolutional Neural Networks (CNNs)28, 29, 30 and later, Graph Neural Networks (GNNs),31, 32, 33, 34, 35, 36 made great gains in their application to molecular property prediction, drug discovery,37 and synthesis prediction.38 Such methods were able to capture complex patterns in data, and therefore enabled the identification of novel materials for high-impact needs such as energy storage and conversion.39, 40
In this review, we explore the next phase of AI in chemistry, namely the use of Large Language Models (LLMs) and autonomous agents. Inspired by successes in natural language processing (NLP), LLMs were adapted for chemical language (e.g., SMILES41) to tackle tasks from synthesis prediction to molecule generation42, 43, 44. We will then explore the integration of LLMs into autonomous agents as illustrated by M. Bran et al. 45 and Boiko et al. 46, which may be used for data interpretation or, for example, to experiment with robotic systems. We are at a crossroads where AI enables chemists to solve major global problems faster and streamline routine lab tasks. This enables, for instance, the development of larger, consistent experimental datasets and shorter lead times for drug and material commercialization. As such, language has been the preferred mechanism for describing and disseminating research results and protocols in chemistry for hundreds of years.47
The review structure is as follows: Section 1.1 discusses current challenges that need to be addressed to increase the impact of AI in chemistry. Section 2 introduces transformers, covering encoder-only, decoder-only, and encoder-decoder architectures. Section 3 surveys work with LLMs, linking each transformer architecture to suitable chemistry areas. Section 4 describes autonomous agents and their applications in chemistry research. Section 5 surveys LLM-based agents in chemistry. Section 6 discusses future challenges and opportunities, followed by the conclusion in Section 7. We distinguish between “text-based” and “mol-based” inputs and outputs, with “text” referring to natural language and “mol” referring to the chemical syntax for material structures, as introduced by Zhang et al. 48.
1.1 Challenges in chemistry
With the potential to accelerate scientific discovery through the use of AI, it makes sense to connect this potential to some of the more pressing opportunities in the chemical sciences as may be categorized under three broad umbrellas: Property Prediction, Property-Directed Molecule Generation, and Synthesis Prediction. The first challenge is to predict a property for a given compound to decide if it should be synthesized for a specific application, such as an indicator,49 light harvester,50 or catalyst.51 To have better models for property prediction, good quality data is crucial. We discuss the caveats and issues with the current datasets in Section 3.1 and illustrate state-of-the-art findings in Section 3.2.
The second challenge is to generate novel chemical structures that meet desired chemical profiles or specific properties.52 Success would accelerate progress in various chemical applications, but reliable reverse engineering (inverse design)53 is not yet feasible over the vast chemical space.54 For instance, inverse design coupled with automatic selection of novel structures (de novo molecular design) could develop a drug targeting a specific protein while retaining properties like solubility, toxicity, and blood-brain barrier permeability.55 The complexity of connecting de novo design with property prediction is high and raises ethical questions that need to be addressed to guide the development of such models. We show how state-of-the-art models currently perform in Section 3.3.
The third major challenge is predicting its optimal synthesis using inexpensive, readily available, and non-toxic starting materials. There will always likely be a solution within a large chemical space: an alternative molecule with similar properties that is easier to synthesize. Exploring this space to find a new molecule with the right properties and a high-yield synthesis route unites these challenges. The number of possible stable chemicals is estimated to be up to $10^{180}$.56, 57, 58, 59 Exploring this vast space requires significant acceleration beyond current methods.60 Restrepo 56 emphasizes the need to catalog data on failed syntheses to build a comprehensive dataset of chemical features. Autonomous chemical resources can accelerate database growth and tackle this challenge. Thus, automation is considered a fourth major challenge in chemistry.61, 62, 63, 64 The following discussion explores how LLMs and autonomous agents can provide the most value. Relevant papers are discussed in Section 3.4
2 Large Language Models
The prior state-of-the-art for seq2seq was considered to be the Recurrent Neural Network (RNN),65 typically as implemented in Hochreiter and Schmidhuber 66. The RNN retains “memory” of previous steps in a sequence to predict later parts. However, as sequence length increases, gradients can become vanishingly small or explosively large67, 68, preventing effective use of earlier information in long sequences. RNNs have thus fallen behind Large Language Models (LLMs).
LLMs are deep neural networks (NN) with billions of parameters. Mostly, LLMs implement a transformers architecture, introduced by Vaswani et al. 69, although other architectures for longer input sequences [^1] These longer sequence models are more important in biology, where very long genetic sequences are relevant are being actively explored70, 71, 72. Transformers are well-developed in chemistry and are the dominant paradigm behind nearly all state-of-the-art sequence modeling results and, thus, the focus. More detailed discussion on current LLMs applied to general purposes can be found elsewhere.73
2.1 The Transformer
Figure 1: a) The generalized encoder-decoder transformer: The encoder on the left converts an input into a vector, while the decoder on the right predicts the next token in a sequence. b) Encoder-decoder transformers are traditionally used for translation tasks and, in chemistry, for reaction prediction, translating reactants into products. c) Encoder-only transformers provide a vector output and are typically used for sentiment analysis. In chemistry, they are used for property prediction or classification tasks. d) Decoder-only transformers generate likely next tokens in a sequence. In chemistry, they are used to generate new molecules given an instruction and description of molecules.
The transformer was introduced in, “Attention is all you need” by Vaswani et al. 69 in 2017. A careful line-by-line review of the model can be found in “The Annotated Transformer” 74. The transformer was the first sequence-to-sequence (seq2seq) model based entirely on attention. Here, a seq2seq model processes a text input, such as an English paragraph, and produces a different text sequence, such as a French translation of that paragraph. The concept of “attention” is a focus applied to certain words of the input, which would convey the most importance, or the context of the passage, and thereby would allow for better decision-making and greater accuracy. However, in a practical sense, “attention” is implemented simply as the dot-product between token embeddings with some non-linear learned function, and is described further below.
The transformer architecture implements two main modules: the encoder and the decoder. Figure 1a provides a cartoon diagram of the general encoder-decoder transformer architecture. The input is first tokenized, from the model’s vocabulary,75, 76, 77, 78, 79 into computable integers that are then converted into numerical vectors using embedding layers. A single encoder stack is shown here, where in reality there are many, with six in the original model.69 The later stacks each use outputs from the prior stack, typically as a $\sim$512 per-token vector. Each encoder applies positional encoding typically using sine and cosine functions, with a frequency depending on that word’s position, to handle sequences of any length. Other functions can also be used.80 The encoder stack then comprises a multi-headed self-attention mechanism, which relates each “word” to others in the sequence by computing attention scores based on queries, keys, and values. This is followed by normalization and residual connections. Residual connection is represented in Figure 1a by the “by-passing” arrows, which indicate how each sub-layer output is added to its input, which helps mitigate the vanishing gradient problem.67, 68 Lastly, the output is further refined through a pointwise feed-forward network with an activation function, resulting in a set of vectors representing the input sequence with rich contextual understanding, ready for the decoder.
Similarly, the decoder workflow begins with converting the initial input token into numerical vectors, known as the output embedding step, as this initial input serves as the start token for the model’s final output. Positional encodings are then added, followed by a stack of approximately six identical decoder layers. Each decoder layer includes a Masked Self-Attention Mechanism to prevent positions from attending to subsequent ones and an Encoder-Decoder Multi-Head Attention to align encoder outputs with decoder attention layer outputs (indicated by the square connecting arrows in Figure 1a), focusing on relevant parts of the input. A fully connected Feed-Forward Neural Network refines the output further. The final layer acts as a linear classifier, mapping the output to the vocabulary size, and a softmax layer converts this output into probabilities, with the highest indicating the predicted next word. Each sub-layer is followed by normalization and includes a residual connection. The decoder generates output sequences autoregressively, incorporating each newly predicted token until an end token signals completion. Figure 1b shows a cartoon illustration of the encoder-decoder model’s strength in translating between languages or from atoms in reactants mapped into new positions in associated products. Figure 1c depicts the ability of an encoder-only model to detect an additional meaning, or property, from an input sequence, represented as a number description. Figure 1d presents a cartoon depiction of a decoder-only architecture inferring and generating a consequence, based on an original input. Hence, a new peptide sequence has been inferred after the combined input of two smaller amino acids, indicating how the decoder-only approach can be used to generate potentially new compounds.
2.2 Model training
The common lifetime of an LLM consists of being first pretrained using unsupervised or self-supervised techniques, generating what is called a base model. This base model is then fine-tuned for specific applications using supervised techniques. Finally, the supervised fine-tuned model is further tuned with reward models to improve human preference or some other non-differentiable and sparse desired character81.
Unsupervised pretraining
A significant benefit implied in all the transformer models described in this review is that unsupervised learning takes place with a vast corpus of text. Thus, the algorithm learns patterns from untagged data, which opens up the model to larger datasets that may not have been explicitly annotated by humans. The advantage is to discover underlying structures or distributions without being provided with explicit instructions on what to predict, nor with labels that might indicate the correct answer.
Supervised fine-tuning
After this pretraining, many models described herein are fine-tuned on specific downstream tasks (e.g., text classification, question answering) using supervised learning. In supervised learning, models learn from labeled data, and map inputs to known outputs. Such fine-tuning allows the model to be adjusted with a smaller, task-specific dataset to perform well on that downstream task.
LLM alignment
A key task after model training is aligning its output with what human users prefer. This can be done for improving LLM style and tone. Or, it can be to reduce harmful outputs because the training objectives used during the pretraining and the fine-tuning do not include human values.82 Reinforcement learning (RL) techniques are among the most common alignment techniques. RL is a type of machine learning where actions are performed and feedback is received through rewards, the goal being to learn a policy that maximizes the cumulative reward. However, human preferences can only be evaluated on complete outputs from models, preventing supervised fine-tuning (which requires token-by-token losses).83, 84, 85 To apply RL, sampling tokens from the model is recast as a Markov Decision Process (MDP) - actions are tokens and the reward is zero until the EOS (end of sequence) token, at which point the human preference reward is applied86 Reinforcement Learning with Human Feedback (RLHF)87 is the most used alignment technique. However, alternative strategies include reinforcement learning with synthetic feedback (RLSF)88, proximal policy optimization (PPO)89, REINFORCE90, and others.83, 86, 91, 92. A general discussion about RL potential for LLM fine-tuning was published by Cao et al. 93 and Shen et al. 82.
There are ways to reformulate the RLHF process into a direct optimization problem with a different loss. This is known as direct preference optimization (DPO)94 and a popular competitor to PPO (the most commonly used approach) due to its simplicity. It overcomes the lack of token-by-token loss signal by comparing two completions at a time. The discussions about which technique is superior remain very active in the literature.95
Finally, the alignment may not be to human preferences but to downstream tasks that do not provide token-by-token rewards. For example, Bou et al. 96 and Hayes et al. 97 both use RL on a language model for improving its outputs on a downstream scientific task.
2.3 Model types
While the Vaswani Transformer69 employed an encoder-decoder structure for sequence-to-sequence tasks, the encoder and decoder were ultimately seen as independent models, leading to “encoder-only”, and “decoder-only” models described below.
2.3.1 Encoder-only models
Beyond Vaswani’s transformer,69 used for sequence-to-sequence tasks, another significant evolutionary step forward came in the guise of the Bidirectional Encoder Representations from Transformers, or “BERT”, described in October 2018 by Devlin et al. 98 BERT utilized only the encoder component, achieving state-of-the-art performance on sentence-level and token-level tasks, outperforming prior task-specific architectures.98 The key difference was BERT’s bidirectional transformer pretraining on unlabeled text, meaning the model looks at the context both to the left and right of the word in question, facilitated by a Masked Language Model (MLM). This allows a richer understanding of the sentence. The encoder-only design meant BERT focused on generating a deeper "understanding" of the input sequence, rather than mapping input sequences to output sequences. In pretraining, BERT also uses Next Sentence Prediction (NSP). “Sentence” here means an arbitrary span of contiguous text. The MLM task randomly masks tokens and predicts them by looking at preceding and following contexts simultaneously, inspired by Taylor99. NSP predicts whether one sentence logically follows another, training the model to understand sentence relationships. This bidirectional approach allows BERT to recognize greater nuance and richness in the input data.
Subsequent evolutions of BERT include, for example, RoBERTa, (Robustly optimized BERT approach), described in 2019 by Facebook AI’s Liu et al. 100. RoBERTa was trained on a larger corpus, for more iterations, with larger mini-batches, and longer sequences, improving model understanding and generalization. By removing the NSP task and focusing on the MLM task, performance improved. RoBERTa dynamically changed masked positions during training and used different hyperparameters. Evolutions of BERT also include domain-specific pretraining and creating specialist LLMs for fields like chemistry, as described below (see Section 3).
2.3.2 Decoder-only models
In June 2018, Radford et al. 101 from OpenAI proposed the Generative Pretrained Transformer (GPT) in their paper, “Improving Language Understanding by Generative Pretraining”. GPT used a decoder-only, left-to-right unidirectional language model to predict the next word in a sequence based on previous words, without an encoder. Unlike the Vaswani Transformer’s decoder, GPT could predict the next sequence, applying general language understanding to specific tasks with smaller annotated datasets.
GPT employed positional encodings to maintain word order in its predictions. Unlike the Vaswani Transformer, GPT’s self-attention mechanism prevented tokens from attending to future tokens, ensuring each word prediction depended only on preceding words. Hence a decoder-only architecture represents a causal language model, one that generates each item in a sequence based on the previous items; the generation of each subsequent output is causally linked to the history of generated outputs and nothing ahead of the current word affects its generation.
2.3.3 Encoder-decoder models
Evolving further, BART (Bidirectional and Auto-Regressive Transformers) was introduced by Lewis et al. from Facebook AI in 2019.102 BART combined the context learning strengths of the bidirectional BERT, and the autoregressive capabilities of models like GPT, which excel at generating coherent text. BART was thus a hybrid seq2seq model, consisting of a BERT-like bidirectional encoder and a GPT-like autoregressive decoder. This is nearly the same architecture as Vaswani et al. 69; the differences are in the pretraining. BART was pretrained using a task that corrupted text by, for example, deleting tokens, and shuffling sentences. It then learned to reconstruct the original text with left-to-right autoregressive decoding as in GPT models.
2.3.4 Multi-task and multi-modal models
In previous sections, we discussed LLMs that take natural language text as input and then they output either a learned representation or another text sequence. These models traditionally perform tasks like translation, summarization, and classification. Raffel et al. 103 developed the Text-to-Text Transfer Transformer (T5) to demonstrate that various tasks can be reframed as text-to-text tasks, allowing the same model architecture and training procedure to be used with the same set of weights across different tasks. This approach has inspired the development of multi-task models that can adapt to different tasks at inference time. For example, Flan-T5104 used instruction fine-tuning with chain-of-thought prompts, enabling it to generalize to unseen tasks, such as generating rationales before answering. More complex approaches have been proposed for robust multi-task models.105, 106, 107, 108
Additionally, LLMs have been extended to understand different input modalities, despite initially receiving only text. For instance, Fuyu109 uses linear projection to adapt image representations into the token space of an LLM, enabling a decoder-only model to write figure captions. Scaling up, next-GPT110 was developed as an “any-to-any” model, processing multiple modalities—text, audio, image, and video—through modality-specific encoders. The encoded representation is projected into a decoder-only token space, and the LLM’s output is processed by a domain-specific diffusion model to generate each modality’s output. Multitask or multimodel methods are further described below as these methods start to connect LLMs with autonomous agents.
3 LLMs for chemistry and biochemistry
Integrating LLMs into chemistry and biochemistry has revolutionized molecular design, synthesis prediction, and property analysis. These models, trained on vast datasets, interpret chemical languages like SMILES and InChI to predict molecular behavior accurately. This section explores the importance of trustworthy datasets and the necessity of good benchmarks and LLM applications in molecular representations, property prediction, inverse design, and synthesis prediction. Figure 3 illustrates the capabilities of different LLMs available currently, and Figure 4 shows a chronological map of LLMs in chemistry and biology.
3.1 Molecular Representations, Datasets, and Benchmarks
Given the diversity of data in chemistry-based machine learning, and the many different ways that a compound can be described, from molecular structure to property description, there are multiple ways a molecule can be represented, underlining this heterogeneity.111, 112, 113, 114 Some common forms include molecular graphs115, 116, 117, 3D point clouds118, 119, 120, 121, and quantitative feature descriptors.122, 123, 124, 125, 112 In this review, we will focus on string-based representations of molecules, given the interest in language models. Among the known string representations, we can cite IUPAC names, SMILES,41 DeepSMILES,126 SELFIES,127 and InChI,128 as recently reviewed by Das et al. 129.
Figure 2: Number of training tokens (on log scale) available from various chemical sources compared with typical LLM training runs. The numbers are drawn from ZINC130, PubChem131, Touvron et al. 132, ChEMBL133, and Kinney et al. 134
Regarding datasets, there are two types of data used for training LLMs: training data and evaluation data. Training data should be grounded on real molecules to build the correct “prior belief” of a molecule definition during the LLM training process. Similar care is taken when natural language training data used in GPT-4, for instance, is built from real sentences or code, not random combinations of words that might be grammatically correct. Figure 2 shows the number of tokens on common chemistry datasets in comparison with the LLaMA2 training data, extracted from References [130, 131, 132, 133, 134]. According to this data, the largest chemical training corpus, composed of hypothetical chemical structures, composes billions of tokens against a few trillion in the LLaMA 2 training data. Interestingly, Segler et al. 135 demonstrated that even using the Reaxys dataset, a very small human-curated set of chemical reactions, is enough to enable state-of-the-art retrosynthesis results. Therefore, the lack of data is not the only problem. Instead, the lack of good-quality data seems to be the pivotal factor holding back the development of better scientific LLMs. Figure 2 illustrates the number of tokens on ZINC, a dataset with hypothetical chemical structures130
On the other hand, benchmark data evaluates the success of these models. Numerous datasets, curated by the scientific community, are available for this purpose136, 137. Among them, MoleculeNet55, first published in 2017, is the most commonly used labeled dataset for chemistry. However, MoleculeNet has several issues: it is small, contains errors and inconsistencies, and lacks relevance to real-world chemistry problems138, 139, 140, 141. Pat Walters, a leader in ML for drug discovery, stated, “I think the best way to make progress on applications of machine learning to drug discovery is to fund a large public effort that will generate high-quality data and make this data available to the community”142.
Walters makes several constructive critiques such as, for example, how the QM7, QM8, and QM9 datasets, intended for predicting quantum properties from 3D structures, are often misused with predictions based incorrectly on 1D SMILES strings, which inadequately represent 3D molecular conformations. Walters also makes suggestions for more relevant benchmarks, and also datasets with more valid entries. He suggests the Absorption, Distribution, Metabolism, and Excretion (ADME) data curated by Fang et al. 143, and the Therapeutic Data Commons (TDC)144, 145 and TDC-2146. These datasets contain measurements of real compounds, making them grounded in reality. Moreover, ADME is crucial for determining a drug candidate’s success, while therapeutic results in diverse modalities align with metrics used in drug development.
Here, we hypothesize that the lack of easily accessible, high-quality data in the correct format for training foundational chemical language models is a major bottleneck to the development of the highly desired “super-human” AI-powered digital chemist. In addition, we argue that extracting data from scientific chemistry papers might be an interesting commitment to generating data of high quality, grounded to the truth, and on a large scale.
{forest}
Figure 3: Classification of LLMs in chemistry and biochemistry according to their application.
Figure 4: Illustration of how Large Language Models (LLMs) evolved chronologically.
3.2 Property Prediction and Encoder-only mol-LLMs
Table 1: Encoder-only scientific LLMs. aa:“Model Size” is reported as the number of parameters. $b$: The authors report they not used as many encoder layers as it was used in the original BERT paper. But the total number of parameters was not reported.
LLM Model Sizea Training Data Architecture Application Release date
CatBERTa164 355M OpenCatalyst2020 (OC20) RoBERTa Property prediction 2023.11
SELFormer176 $\sim$86M $\sim$2M compounds from ChEMBL Property prediction RoBERTa 2023.05
MaterialsBERT184 110M 2.4M material science abstracts + 750 annotated abstract for NER BERT NER and property extraction 2023.04
SolvBERT185 b 1M SMILES of solute-solvent pairs from CombiSolv-QM and LogS from Boobier et al. 186 BERT Property prediction 2023.01
MolFormer187 $b$ PubChem and ZINC BERT Property prediction 2022.12
ChemBERTa-2177 5M - 46M 77M SMILES from PubChem RoBERTa Property prediction 2022.09
ScholarBERT165 340M, 770M Public.Resource.Org, Inc BERT Property prediction 2022.05
MatSciBERT160 110M $\sim$150K material science paper downloaded from Elsevier BERT NER and text classification 2022.05
MatBERT168 110M Abstracts from solid state articles and abstracts and methods from gold nanoparticle articles BERT NER 2022.04
PubMedBERT161 110M 14M abstracts from PubMed BERT NER, QA, and document classification 2021.10
Mol-BERT179 110M $\sim$4B SMILES from ZINC15 and ChEMBL27 BERT Property prediction 2021.09
ChemBERT169 110M $\sim$200k extracted using ChemDataExtractor BERT NER 2021.06
MolBERT181 $\sim$85M ChemBench BERT Property prediction 2020.11
ChemBERTa42 10M SMILES from PubChem RoBERTa Property prediction 2020.10
BioMegatron170 345M, 800M, 1.2B Wikipedia, CC-Stories, Real-News, and OpenWebtext Megatron-LM NER and QA 2020-10
BioBERT171 110M PubMed and PMC BERT NER and QA 2020.02
SciBERT162 110M 1.14M papers from Semantic Scholar BERT NER and sentence classification 2019.09
SMILES-BERT188 $b$ $\sim$18M SMILES from ZINC BERT Property prediction 2019.09
BlueBERT163 110M PubMed and MIMIC-III BERT NER, and document classification 2019.06
ClinicalBERT166 110M MIMIC-III BERT Pacient readmission probability 2019.04
Encoder-only transformer architectures solely comprise an encoder. Chemists can leverage the simpler encoder-only architecture for tasks where the primary goal is classification, efficient exploration of chemical space, and property prediction. Since encoder-only architectures are mostly applied to property prediction, we describe here the relative importance of this principal chemical challenge. Sultan et al. 189 also discussed the high importance of this task.
3.2.1 Property Prediction
The universal value of chemistry lies in identifying and understanding the properties of compounds to optimize their practical applications. In the pharmaceutical industry, therapeutic molecules interact with the body in profound ways.190, 191, 192 Understanding these interactions and modifying molecular structures to enhance those therapeutic benefits can lead to significant medical advancements.193 Similarly, in polymer science, material properties depend on chemical structure, polymer chain length, and packing,194 and a protein’s function similarly depends on its structure and folding. Historically, chemists have identified new molecules from natural products195 and screened them against potential targets196 to test their properties for diseases. Once a natural product shows potential, chemists synthesize scaled-up quantities for further testing or derivatization,197, 198, 199 a costly and labor-intensive process.200, 201 Traditionally, chemists have used their expertise to hypothesize the properties of new molecules derived from those natural products, hence aiming for the best investment of synthesis time and labor. To support the chemical industry in more accurate property prediction, computational chemistry has evolved.202 Quantum theoretical calculations predict certain properties reliably, while force-field-based Molecular Dynamics (MD)203 predict packing and crystal structures of large molecular ensembles, though both require substantial computational resources.204, 205, 206, 207, 208 Property prediction can now be enhanced through machine learning tools,209, 124, 210, 211 and more recent advancements in LLMs lead to effective property prediction without the extensive computational demands of quantum mechanics and MD calculations. Combined with human insight, AI can revolutionize material development, enabling the synthesis of new materials with a high likelihood of possessing desired properties for specific applications.
3.2.2 Encoder-only Mol-LLMs
Encoder-only models are exemplified by the BERT architecture, which is commonly applied in natural language sentiment analysis to extract deeper patterns from prose.212 The human chemist has been taught to look at a 2D image of a molecular structure and to recognize its chemical properties or classify the compound. Therefore, encoder-only models would ideally convert SMILES strings, empty of inherent chemical essence, into a vector representation, or latent space, reflecting those chemical properties. This vector representation can then be used directly for various downstream tasks.
Schwaller et al. 213 used a BERT model to more accurately classify complex synthesis reactions by generating reaction fingerprints from raw SMILES strings, without the need to separate reactants from reagents in the input data, thereby simplifying data preparation. The BERT model achieved higher accuracy (98.2%) compared to the encoder-decoder model (95.2%) for classifying reactions. Accurate classification aids in understanding reaction mechanisms, vital for reaction design, optimization, and retrosynthesis. Toniato et al. 214 also used a BERT architecture to classify reaction types for downstream retrosynthesis tasks that enable manufacture of any molecular target. Further examples of BERT use include unsupervised reaction atom-to-atom mapping 215, 216. These chemical classifications would accelerate research and development in organic synthesis, described further below.
Due to limited labeled data for molecular property prediction, Wang et al. proposed a semi-supervised SMILES-BERT model, pretrained on a large unlabeled dataset with a Masked SMILES Recovery task.182 The model was then fine-tuned for various molecular property prediction tasks, outperforming state-of-the-art methods in 2019 on three chosen datasets varying in size and property. This marked a shift from using BERT for reaction classification towards property prediction and drug discovery.
Zhang et al. 217 also tackled the issue of limited property-labeled data and the lack of correlation between any two datasets labeled for different properties, hindering generalizability. They introduced multitask learning BERT (MTL-BERT), which used large-scale pretraining and multitask learning with unlabeled SMILES strings from ChEMBL.133 This approach mined contextual information and extracted key patterns from complex SMILES strings, improving model interpretability. The model was fine-tuned for relevant downstream tasks, achieving better performance than state-of-the-art methods in 2022 on 60 molecular datasets from ADMETlab218 and MoleculeNet.55
In 2021, Li and Jiang179 introduced Mol-BERT, pretrained on four million unlabeled drug SMILES from the ZINC15219 and ChEMBL27133 databases to capture molecular substructure information for property prediction. Their work leveraged the underutilized potential of large unlabeled datasets like ZINC. Mol-BERT consisted of three components: PretrainingExtractor, Pretraining Mol-BERT, and Fine-Tuning Mol-BERT. It treated Morgan fingerprint fragments as “words” and compounds as “sentences,” using RDKit and the Morgan algorithm for canonicalization and substructure identification. This approach generated comprehensive molecular fingerprints from SMILES strings, used in a Masked Language Model (MLM) task for pretraining. Mol-BERT was fine-tuned on labeled samples, providing outputs as binary values or continuous scores for classification or regression, and it outperformed existing sequence and graph-based methods by at least 2% in ROC-AUC scores on Tox21, SIDER, and ClinTox datasets.55
Ross et al. 178 introduced MoLFormer, a large-scale unsupervised BERT model for more accurate and faster molecular property predictions than Density Functional Theory calculations or wet-lab experiments. They trained MoLFormer with rotary positional embeddings on SMILES sequences of 1.1 billion unlabeled molecules from PubChem131 and ZINC219. Rotary positional encoding captures token positions more effectively than traditional methods, 69 improving modeling of sequence relationships. MoLFormer outperformed state-of-the-art GNNs on several classification and regression tasks from ten MoleculeNet55 datasets, while performing competitively on two others. It effectively learned spatial relationships between atoms, predicting various molecular properties, including quantum-chemical properties.
With ChemBERTa, Chithrananda et al. 42 explored the impact of pretraining dataset size, tokenization strategy, and the use of SMILES or SELFIES, distinguishing their work from other BERT studies. They used HuggingFace’s RoBERTa transformer,220 and referenced a DeepChem55 tutorial for accessibility. Their results showed improved performance on downstream tasks (BBBP, ClinTox, HIV, Tox21 from MoleculeNet55) as the pretraining dataset size increased from 100K to 10M. Although ChemBERTa did not surpass state-of-the-art GNN-based baselines like Chemprop (which used 2048-bit Morgan Fingerprints from rdkit),221 the authors suggested that with expansion to larger datasets they would eventually beat those baselines. The authors compared Byte-Pair Encoder (BPE) with a custom SmilesTokenizer and its regex developed by Schwaller et al.222 while exploring tokenization strategies. They found the SmilesTokenizer slightly outperformed BPE, suggesting more relevant sub-word tokenization is beneficial. No difference was found between SMILES and SELFIES, but the paper highlighted how attention heads in transformers could be visualized with BertViz,223 showing certain neurons selective for functional groups. This study underscored the importance of appropriate benchmarking and addresses the carbon footprint of AI in molecular property prediction.
In ChemBERTa-2, Ahmad et al. 177 aimed to create a foundational model applicable across various tasks. They addressed criticisms of biased data in LLMs by training on 77M samples and adding a Multi-Task Regression component to the pretraining. ChemBERTa-2 matched state-of-the-art architectures on MoleculeNet.55 As with ChemBERTa, the work was valuable because of additional exploration, in this case into how pretraining improvements affected certain downstream tasks more than others, depending on the type of fine-tuning task, the structural features of the molecules in the fine-tuning task data set, or the size of that fine-tuning dataset. The result was that pretraining the encoder-only model is important, but gains could be made by considering the chemical application itself, and the associated fine-tuning dataset.
In June 2023, Yuksel et al. 176 introduced SELFormer, building on ideas from ChemBERTa2177 and using SELFIES for large data input. Yuksel et al. 176 argue that SMILES strings have validity and robustness issues, hindering effective chemical interpretation of the data. SELFormer uses SELFIES and is pretrained on two million drug-like compounds, fine-tuned for diverse molecular property prediction tasks (BBBP, SIDER, Tox21, HIV, BACE, FreeSolv, ESOL, PDBbind from MoleculeNet).55 SELFormer outperformed all competing methods for some tasks and produced comparable results for the rest. It could also discriminate molecules with different structural properties. The paper suggests future directions in multimodal models combining structural data with other types of molecular information, including text-based annotations. We will discuss such multimodal models below.
Also in 2023, Yu et al. 185 published SolvBERT, a multi-task BERT-based regression model that could predict both solvation free energy and solubility from the SMILES notations of solute-solvent complexes. It was trained on the CombiSolv-QM dataset,224 a curation of experimental solvent free energy data called CombiSolv-Exp-8780,225, 226, 227, 228 and the solubility dataset from Boobier et al. 186. SolvBERT’s performance was benchmarked against advanced graph-based models229, 230 This work is powerful because there is an expectation that solvation free energy depends on 3-dimensional conformational properties of the molecules, or at least 2D properties that would be well characterized by graph-based molecular representations. SolvBERT showed comparable performance to DMPNN in predicting solvation free energy, largely due to its effective clustering feature in the pretraining phase as shown by TMAP visualizations. Furthermore, SolvBERT outperformed GROVER in predicting experimentally evaluated solubility data for new solute-solvent combinations. This underscores the significance of SolvBERT’s ability to capture the dynamic and spatial complexities of solvation interactions in a text-based model.
Despite significant progress, Jiang et al. 231 highlighted the limitations of labeled data in 2024 and introduced INTransformer,231 a transformer-based method for predicting molecule properties. INTransformer enhances the capture of global molecular information by adding perturbing noise and employing contrastive learning to artificially augment smaller datasets. This method achieved higher performance, and ongoing work continues to explore various transformer strategies for smaller datasets.
3.3 Property Directed Inverse Design and Decoder-only mol-LLMs
Table 2: Decoder-only scientific LLMs. aa:“Model Size” is reported as the number of parameters. “PubMed” refer to the PubMed abstracts dataset, while PMC (PubMed Corpus) refers to the full-text corpus dataset. $b$: The total number of parameters was not reported.
LLM Model Sizea Training Data Architecture Application Release date
Tx-LLM232 $b$ TDC datasets PaLM-2 Property prediction and retrosynthesis 2024.06
BioMedLM147 2.7B PubMed abstracts and full articles GPT QA 2024.03
BioMistral148 7B PubMed Central (PMC) Mistral QA 2024.02
BiMediX149 8x7B 1.3M Arabic-English instructions (BiMed) Mixtral QA 2024.02
EpilepsyLLM150 7B Data from the Japan Epilepsy Association, Epilepsy Information Center, and Tenkan Net LLaMa QA 2024.01
ChemSpaceAL233 $b$ ChEMBL 33, GuacaMol v1, MOSES, and BindingDB 08-2023 GPT Molecule Generation 2024.01
CheXagent151 7B 28 publicly available datasets, including PMC, MIMIC, wikipedia, PadChest, and BIMCV-COVID-19 Mistral QA, Image understanding 2024.01
BioMedGPT-LM152 7B and 10B 5.5M biomedical papers from S2ORC LLaMA2 QA 2023.08
Darwin153 7B SciQ and Web of Science LLaMA QA, Property prediction, NER, and Molecule Generation 2023.08
cMolGPT44 $b$ MOSES GPT Molecule Generation 2023.04
PMC-LLaMA154 7B and 13B MedC-k and MedC-I LLaMA QA 2023.04
Regression Transformer173 $\sim$27M ChEMBL, MoleculeNet, USPTO, etc XLNet Property prediction, Molecule tunning, Molecule generation 2023.04
Galactica155 125M, 1.3B, 6.7B, 30B, 120B The galactica corpus, a curation with 62B scientific documents Decoder-only QA, NER, Document Summarization, Property Prediction 2022.11
BioGPT156 355M 15M of Title and abstract from PubMed GPT-2 QA, NER, and Document Classification 2022-09
MolGPT234 6M MOSES and GuacaMol GPT Molecule Generation 2021.10
Adilov2021235 13.4M 5M SMILES from ChemBERTa’s PubChem-10M. GPT-2 Property prediction and molecule generation 2021.09
Decoder-only GPT-like architectures offer significant value for property directed molecule generation and de novo chemistry applications because they excel in generating novel molecular structures by learning from vast datasets of chemical compounds, and capture intricate patterns and relationships within molecular sequences. They thus propose viable new compounds that adhere to desired chemical properties and constraints, enabling rapid exploration and innovation within an almost infinite chemical space. We first describe property directed inverse design from a chemistry perspective, and then go on to provide examples of how decoder-only LLMs have propelled inverse design forward.
3.3.1 Property Directed Inverse Design
Nature has been a significant source of molecules that inhibit disease proliferation, as organisms have evolved chemicals to protect themselves. This has led to most pharmaceuticals being derived from natural products,236, 237 which offer advantages such as cell permeability, target specificity, and a vast chemical diversity.238 However, despite these benefits, there are high costs and complexities associated with high-throughput screening and synthesis of natural products.238, 236 While building on the chemical diversity and efficacy of natural products, AI also opens new avenues for synthesizing unique compounds efficiently. Advanced in-silico molecular design allows for rapid mutation239 towards valid de-novo molecular structures that are synthesizable,211, 240 streamlining the iterative process of drug development. Yet, the true innovation lies in the potential of LLM-driven target-directed molecular design or “Inverse Design,” where we can start with a desired property and directly generate molecules that manifest this attribute, bypassing traditional stepwise modifications.241 This capability dramatically accelerates the pathway from concept to viable therapeutic agents and aligns well with decoder-only LLM architectures.
3.3.2 Decoder-only mol-LLMs
In 2021, Adilov 235 presented “Generative pretraining from Molecules,” one of the first applications of decoder-only models to SMILES strings. It pretrained a GPT-2-like causal transformer for self-supervised learning, introducing “adapters” between attention blocks for task-specific fine-tuning.242 This method, requiring minimal architectural changes, offered versatility in molecule generation and property prediction, aiming to surpass ChemBERTa’s encoder-only performance with a more scalable and resource-efficient approach. Another early decoder-only SMILES-based model was Bagal et al’s MolGPT model,234. MolGPT, with a mere 6 million parameters, advanced GPT-type LLMs for molecular generation. Its decoder-only architecture with masked self-attention facilitated learning long-range dependencies, enabling chemically valid SMILES representations that met complex structural rules involving valency and ring closures. The paper also used salience measures for interpretability in predicting SMILES tokens. MolGPT outperformed many existing Variational Auto-Encoder (VAE) based approaches243, 244, 245, 246, 247, 248, 249, 250 in predicting novel molecules with specified properties, being trained on datasets like MOSES251 and GuacaMol.252 It showed good performance with metrics like validity, uniqueness, Frechet ChemNet Distance (FCD),253 and KL divergence,252. While MolGPT’s computational demands might be higher than traditional VAEs, its ability to generate high-quality, novel molecules justifies this trade-off, but future research could likely optimize model efficiency or explore lighter versions. A brief summary of advancements in transformer-based models for de-novo molecule generation from 2023 and 2024 follows.
Haroon et al. 254 further developed a GPT-based model with relative attention for de novo drug design, showing improved validity, uniqueness, and novelty. Both Wang et al. 255 and Mao et al. 256 presented work that surpassed MolGPT, with Mao et al.’s T5-type model257 generating novel compounds using IUPAC names directly.258 Although T5-based, we include this here because of its relevance to de novo drug design. In a similar vein, Zhang et al. 259 proposed including target 3D structural information in molecular generative models, even though their approach is not LLM-based. However, we nonetheless note its value for future structure-based LLM drug design. They demonstrated that integrating additional biological data significantly enhances the relevance and specificity of generated molecules for targeted drug discovery. Wang et al. 260 discussed PETrans, a deep learning method for generating target-specific ligands using protein-specific encoding and transfer learning. This study emphasized the application of transformer models for generating molecules with a high binding affinity to specific protein targets.
In 2024, Yoshikai et al. 261 discussed the limitations of transformer architectures in recognizing chirality from SMILES representations, highlighting the challenges in learning overall molecular structures with, in particular, chirality impacting the prediction accuracy of molecular properties. They coupled a transformer with a VAE to address this. Qian et al. 262 introduced CONSMI, motivated by contrastive learning in NLP, to generate new molecules using multiple SMILES representations, improving molecular novelty and validity. Kyro et al. 233 presented ChemSpaceAL, an active learning method for protein-specific molecular generation, efficiently discovering molecules with desired characteristics without prior knowledge of existing inhibitors. Yan et al. 263 proposed the GMIA framework, featuring a graph mutual interaction attention decoder for drug-drug interaction prediction, enhancing prediction accuracy and interpretability. Lastly, Shen et al. 264 reported on AutoMolDesigner, an AI-based open-source software for automated design of small-molecule antibiotics.
For a deeper dive into decoder-only transformer architecture in chemistry, we highlight the May 2023 “Taiga” model by Mazuz et al. 85, and cMolGPT by Wang et al. 44. Taiga first learns to map SMILES strings to a vector space, and then refines that space using a smaller dataset of labeled molecules to produce molecules with targeted attributes. It employs an autoregressive mechanism, sequentially predicting each SMILES character based on the previous ones. For optimizing molecular properties, Taiga uses the REINFORCE algorithm90, which develops a strategy to enhance desired molecular features. Although this fine-tuning with reinforcement learning (RL) slightly reduces molecule validity, it significantly boosts their practical applicability. Initially evaluated using the Quantitative Estimate of Drug-likeness (QED) metric,265 Taiga has also demonstrated promising results in targeting IC50 values,133 the BACE protein,266 and anti-cancer activities they collected from a variety of sources. This work underscores the importance of using new models to address applications that require a higher level of chemical sophistication, to illustrate how such models could ultimately be applied outside of the available benchmark datasets. It also builds on the necessary use of standardized datasets and train-validation-test splitting, to demonstrate progress, as explained by Wu et al. 55. Yet, even the MoleculeNet benchmarks55 are flawed and we point the reader here to a more detailed discussion on benchmarking,142 given that a significant portion of molecules in the BACE dataset have undefined stereo centers, which, at a deeper level, complicates the modeling and prediction accuracy.
Wang et al. 44 introduced “cMolGPT: A Conditional Generative Pre-Trained Transformer for Target-Specific de novo Molecular generation,” which underscores the importance of incorporating chemical domain knowledge to effectively navigate the vast landscape of drug-like molecules. Using unsupervised learning and an auto-regressive approach, cMolGPT generates SMILES guided by predefined conditions based on target proteins and binding molecules. Initially trained on the MOSES dataset251 without target information, the model is fine-tuned with embeddings of protein-binder pairs, focusing on generating compound libraries and target-specific molecules for the EGFR, HTR1A, and S1PR1 protein datasets.267, 268, 269, 270
Their approach employs a QSAR model5 to predict the activity of generated compounds, achieving a Pearson correlation coefficient over 0.75. However, reliance on this QSAR model, which has its own limitations, underscores the necessity for more extensive experimental datasets. Despite its capabilities, the cMolGPT model tends to generate molecules within the confines of the same sub-chemical space as the original dataset, indicating that while it successfully identifies potential binders, it faces challenges in broadly exploring the chemical space to discover novel solutions. This limitation highlights the difficulty of generating molecules from significantly different chemical families that are effective in binding or possess a combination of desired properties, suggesting a need for approaches that can more effectively expand the exploration of chemical diversity. Regardless, models capable of this kind of conditional generation are typically trained on large datasets of known protein-ligand interactions. Through this training, they learn the complex relationship between protein structures and the chemical structures of binding ligands. The model learns to predict which molecular features or structures are important for interacting with proteins that have certain characteristics (as captured in the embeddings). While both “Taiga” and “cMolGPT” use transformer architectures and decoder-only setups for molecule generation, they differ in approach. Taiga integrates reinforcement learning to optimize its generative model, while cMolGPT uses target-specific embeddings to guide the generation process.
Outside of training on SMILES strings, Mao et al. 258 introduce iupacGPT, which also uses a decoder-only architecture, modeled after GPT-2. It addresses the limitations of SMILES strings by using IUPAC names to integrate human-readable chemical semantics. It focuses on pretraining with a vast IUPAC dataset and fine-tuning with lightweight networks, excelling in molecule generation, classification, and regression tasks.
We also refer the reader to several other good reviews, 271, 272, 273, 129 with Goel et al. 274 exploring the efficiency of modern machine learning methods in sampling drug-like chemical space for virtual screening and molecular design. The work of Goel et al. 274 discussed generative models, including LLM approaches, for approximating the entire drug-like chemical space and highlighted models conditioned on specific properties or receptor structures.
We provide a segue from this section by introducing the work by Jablonka et al. 275, which showcases a decoder-only GPT model that, despite its training on natural language rather than specialized chemical languages, competes effectively with decoder-only LLMs tailored to chemical languages. The authors use GPT-3 to answer complex chemistry questions and therefore highlight its potential as a foundational tool in the field. This sets the stage for integrating a natural language decoder-only LLM, like ChatGPT, into chemical research, where it could serve as a central hub. This foreshadows future developments to pair the LLM with specialized tools to enhance capabilities, paving the way for the creation of autonomous agents that leverage deep language understanding in scientific domains.
3.4 Synthesis Prediction and Encoder-decoder Mol-LLMs
Table 3: Encoder-decoder scientific LLMs. aa:“Model Size” is reported as the number of parameters. $b$: The total number of parameters was not reported.
LLM Model Sizea Training Data Architecture Application Release date
MOLGEN276 $b$ ZINC15 BART Molecule Generation 2024.03
BioT5+157 252M ZINC20, UniRef50, 33M PubMed articles, 339K mol-text pairs from PubChem, 569K FASTA-text pairs from Swiss-prot T5 Molecule Captioning, Molecule Generation, Property Prediction, 2024-02
BioT5172 252M ZINC20, UniRef50, full-articles from BioRxiv and PubMed, mol-text-IUPAC information from PubChem T5 Molecule Captioning, Property Prediction 2023-12
nach0141 250M MoleculeNet, USPTO, ZINC T5 Property prediction, Molecule generation, Question answering, NER 2023.11
ReactionT5174 220M ZINC and ORD T5 Property prediction and Reaction prediction 2023.11
Text+Chem T5158 60M, 220M 11.5M or 33.5M samples curated from Vaucher et al. 277, Toniato et al. 214, and CheBI-20 T5 Molecule Captioning, Product Prediction, Retrosynthesis, Molecule Generation 2023.01
MolT5159 60M, 770M C4 dataset T5 Molecule Captioning and Molecule Generation 2022.11
T5Chem175 220M USPTO T5 Product Prediction, Retrosynthesis, Property Prediction 2022.03
ChemFormer139 45M, 230M 100M SMILES from ZINC-15 BART Product Prediction, Property Prediction, Molecular Generation 2022.01
SMILES transformer183 $b$ ChEMBL24 Transformer Property prediction 2019.11
Molecular Transformer222 12M USPTO Transformer Product prediction 2019.08
The encoder-decoder architecture is designed for tasks involving the translation of one sequence into another, making it ideal for predicting chemical reaction outcomes or generating synthesis pathways from given reactants. We begin with a background on optimal synthesis prediction and describe how earlier machine learning has approached this challenge. Following that, we explain how LLMs have enhanced chemical synthesis prediction and optimization.
3.4.1 Synthesis Prediction
Once a molecule has been identified through property-directed inverse design, the next challenge is to predict its optimal synthesis, including yield. Shenvi 278 describe how the demanding and elegant syntheses of natural products has contributed greatly to organic chemistry. However, in the past 20 years, the focus has shifted away from complex natural product synthesis towards developing reactions applicable for a broader range of compounds, especially in reaction catalysis.278 Yet, complex synthesis is becoming relevant again as it can be digitally encoded, mined by LLMs, and applied to new challenges. Unlike property prediction, reaction prediction is challenging due to the involvement of multiple molecules. Modifying one reactant requires adjusting all others, with different synthesis mechanisms or conditions likely involved. Higher-level challenges exist for catalytic reactions and complex natural product synthesis. Synthesis can be approached in two ways. Forward synthesis involves building complex target molecules from simple, readily available substances, planning the steps progressively. Retrosynthesis, introduced by E.J. Corey in 1988,279 is more common. It involves working backward from the target molecule, breaking it into smaller fragments whose re-connection is most effective. Chemists choose small, inexpensive, and readily available starting materials to achieve the greatest yield and cost-effectiveness. For example, the first total synthesis of discodermolide280 involved 36 such steps, a 24-step longest linear sequence, and a 3.2% yield. There are many possible combinations for the total synthesis of a target molecule, and the synthetic chemist must choose the most sensible approach based on their expertise and knowledge. However, this approach to total synthesis takes many years. LLMs can now transform synthesis such that structure-activity relationship predictions can be coupled in lock-step with molecule selection based on easier synthetic routes. This third challenge of predicting the optimal synthesis can also lead to the creation of innovative, non-natural compounds, chosen because of such an easier predicted sytnthesis but for which the properties are still predicted to meet the needs of the application. Thus, these three challenges introduced above are interconnected.
3.4.2 Encoder-decoder mol-LLMs
Before we focus on transformer use, some description is provided on the evolution from RNN and Gated Recurrent Unit (GRU) approaches in concert with the move from template-based to semi-template-based to template-free models. Nam and Kim 281 pioneered forward synthesis prediction using a GRU-based translation model, while Liu et al. 282 reported retro-synthesis prediction with a Long Short-Term Memory (LSTM) based seq2seq model incorporating an attention mechanism, achieving 37.4% accuracy on the USPTO-50K dataset.283 The reported accuracies of these early models highlighted the challenges of synthesis prediction, particularly retrosynthesis. Schneider et al. 283 further advanced retrosynthesis by assigning reaction roles to reagents and reactants based on the product.
Evolving from RNNs and GRUs, the field progressed with the introduction of template-based models. Segler and Waller 284 identified that computational “rule-based systems” often failed because they ignored the molecular context, resulting in “reactivity conflicts.” They prioritized suitable transformation rules describing how atoms and bonds change during reactions, applied in reverse for retrosynthesis. They trained a model on 3.5 million reactions, achieving 95% top-10 accuracy in retrosynthesis and 97% for reaction prediction on a validation set of nearly 1 million reactions from the Reaxys database (1771-2015). Though not transformer-based, their work paved the way for LLMs in synthesis applications. However, template-based models rely on explicit reaction templates from known reactions, which limits their ability to predict novel reactions and they require manual updates for learning new data.
Semi-template-based models offered a balance between rigid template-based methods and flexible template-free approaches. They used interpolation or extrapolation within template-defined spaces to predict a wider range of reactions and to adjust based on new data. In 2021, Somnath et al. 285 introduced a graph-based approach recognizing that precursor molecule topology is largely unchanged during reactions. Their model broke the product molecule into “synthons” and added relevant leaving groups, making results more interpretable.286 Training on the USPTO-50k dataset, 283 they achieved a top-1 accuracy of 53.7%, outperforming previous methods.
It is, however, the template-free approaches that align well with transformer-based learning approaches because they learn retrosynthetic rules from raw training data. This provides significant flexibility and generalizability across various types of chemistry. Template-free models are not constrained by template libraries and so can uncover novel synthetic routes that are undocumented or not obvious from existing reaction templates. To pave the way for transformer use in synthesis, Cadeddu et al. 287 drew an analogy between fragments in a compound and words in a sentence due to their similar rank distributions. Schwaller et al. 288 further advanced this with an LSTM network augmented by an attention-mechanism-based encoder-decoder architecture, using the USPTO dataset.283 They introduced a commonly used “regular expression” (or “regex”) for tokenizing molecules, framing synthesis (or retrosynthesis) predictions as translation problems with a data-driven, template-free sequence-to-sequence model. They tracked which starting materials were actual reactants, distinguishing them from other reagents like solvents or catalysts, and used the regex to uniquely tokenize recurring reagents, as their atoms were not mapped to products in the core reaction.
In 2019, Schwaller et al. 222 first applied a transformer for synthesis prediction, framing the task as translating reactants and reagents into the final product. Their model inferred correlations between chemical motifs in reactants, reagents, and products in the dataset (USPTO-MIT,289 USPTO-LEF, 290 USPTO-STEREO288). It required no handcrafted rules and accurately predicted subtle chemical transformations, outperforming all prior algorithms on a common benchmark dataset. The model handled inputs without a reactant-reagent split, following their previous work,288 and accounted for stereochemistry, making it valuable for universal application. Then, in 2020, for automated retrosynthesis, Schwaller et al. 291 developed an advanced Molecular Transformer model with a hyper-graph exploration strategy. The model set a standard for predicting reactants and other entities, evaluated with four new metrics: coverage, class diversity, round-trip accuracy, and Jensen–Shannon divergence. Constructed dynamically, the hypergraph allowed for efficient expansion based on Bayesian-like probability scores, showing high performance despite training data limitations. Notably, accuracy was improved after the re-synthesis of the target product from the newly generated molecular precursors was included in the model. This round-trip accuracy concept was also used by Chen and Jung 292 and Westerlund et al. 293. Also in 2020, Zheng et al. 294 developed a “template-free self-corrected retrosynthesis predictor” (SCROP) using transformer networks and a neural network-based syntax corrector, achieving 59.0% accuracy on a benchmark dataset.295, 283 This approach outperformed other deep learning methods by over 2% and template-based methods by over 6%.
We now highlight advancements in synthesis prediction using the BART Encoder-Decoder architecture, including Chemformer by Irwin et al. 139. This paper emphasized the computational expense of training transformers on SMILES and the importance of pretraining for efficiency. It showed that models pretrained on task-specific datasets or using only the encoder stack were limited for sequence-to-sequence tasks. After transfer learning, Chemformer achieved state-of-the-art results in both sequence-to-sequence synthesis tasks and discriminative tasks, such as optimizing molecular structures for specific properties. They studied the effects of small changes on molecular properties using pairs of molecules from the ChEMBL database133 with a single structural modification. Chemformer’s performance was tested on the ESOL, Lipophilicity, and Free Solvation datasets.55 Irwin et al. 139 also described their use of an in-house property prediction model, but when models train on calculated data for ease of access and uniformity, they abstract away from real-world chemical properties. We again emphasize the importance of incorporating experimentally derived data into Chemistry LLM research to create more robust and relevant models. Continuously curating new, relevant datasets that better represent real-world chemical complexities will enhance the applicability and transferability of these models.
In 2023, Toniato et al. 214 also applied LLMs to single-step retrosynthesis as a translation problem, but increased retrosynthesis prediction diversity by adding classification tokens, or “prompt tokens,” to the target molecule’s language representation, guiding the model towards different disconnection strategies. Increased prediction diversity has high value by providing out-of-the-box synthetic strategies to complement the human chemist’s work. To measure retrosynthesis accuracy, Li et al. 296 introduced Retro-BLEU, a metric adapted from the BLEU score used in machine translation.297 Despite progress in computer-assisted synthesis planning (CASP), not all generated routes are chemically feasible due to steps like protection and deprotection needed for product formation. Widely accepted NLP metrics like BLEU297 and ROUGE298 focus on precision and recall by computing n-gram overlaps between generated and reference texts. Similarly, in retrosynthesis, reactant-product pairs can be treated as overlapping bigrams. Retro-BLEU uses a modified BLEU score, emphasizing precision over recall, as there is no absolute best route for retrosynthesis. Although not yet applied to LLM-based predictions, this approach has value by allowing future performance comparison with a single standard.
Finally, by expanding the use of encoder-decoder architectures outside synthesis prediction into molecular generation, Fang et al. 276 introduced MOLGEN, a BART-based pretrained molecular language model, in a 2023 preprint updated in 2024. MOLGEN addressed three key challenges: generating valid SMILES strings, avoiding an observed bias that existed against natural product-like molecules, and preventing hallucinations of molecules that didn’t retain the intended properties. Pretrained on 100 million molecules using SELFIES127 and a masked language model approach, MOLGEN predicts missing tokens to internalize chemical grammar. An additional highlight of this work is how MOLGEN uses “domain-agnostic molecular prefix tuning.” This technique integrates domain knowledge directly into the model’s attention mechanisms by adding molecule-specific prefixes, trained simultaneously with the main model across various molecular domains. The model’s parameters would thus be adjusted to better capture the complexities and diversities of molecular structures, and domain-specific insights would be seamlessly integrated. To prevent molecular hallucinations, MOLGEN employs a chemical feedback mechanism, to autonomously evaluate generated molecules for appropriate properties, to guide learning and optimization. Such feedback foreshadows a core aspect of autonomous agents, which is their capacity for reflection. We will explore this further below.
3.5 Multi-Modal LLMs
Multitask or multimodal methods applied to chemistry are now described. In 2021, Edwards et al. 299 proposed Text2Mol, which retrieved molecules using natural language descriptions as queries. This task required integrating both molecular and natural language modalities, making it a cross-lingual retrieval problem due to the distinct grammar of molecular representations. The researchers built a paired dataset of molecules with corresponding text descriptions, and developed an aligned semantic embedding space for retrieval. This was enhanced with a cross-modal attention-based model for explainability and reranking. One stated aim was to improve retrieval metrics, which would further advance the ability for machines to learn from chemical literature. In their follow-up work in 2022, Edwards et al. 159 continued using SMILES string representations along with text descriptions. The two key applications were first, to generate captions for molecules from SMILES strings and second, to output molecules based on textual descriptions of desired properties. With this in mind, several important challenges remain. First, molecules can have many possible descriptions based on their diverse properties. For example, aspirin can be described by its pain-relieving effects, its use in preventing heart attacks, its chemical structure (an ester and a carboxylic acid connected to a benzene ring in ortho geometry), or its degradation into salicylic acid and ethanoic acid in moist conditions 300. Describing such a molecule therefore requires expertise in different chemistry domains, depending on its functions and properties. In contrast, image captioning for common objects like cats and dogs needs much less domain expertise. Thus, obtaining a large, paired dataset of chemical representations and textual descriptions is challenging. Another issue is that standard evaluation metrics, like BLEU, are inadequate for molecule-language tasks. To address this, Edwards et al. 159 developed MolT5, which pretrained a model on a large corpus of unlabeled natural language text and molecule strings using a denoising objective. They fine-tuned the model on gold standard annotations and improved the metric based on their prior Text2Mol work 299. MolT5 effectively generated both molecule and caption outputs.
In other work, Seidl et al. 301 developed CLAMP. Trained on large biochemical datasets, CLAMP adapts to new tasks using separate modules for chemical and language inputs, and predicts biochemical activity for drug discovery. Xu et al. 302 presented BioTranslator, a method for translating user-written text descriptions into non-text biological data, facilitating the discovery of novel cell types and applications in protein function prediction and drug target identification. This method enabled data exploration beyond controlled vocabularies using free text. ChatDrug, by Liu et al. 303, integrates multimodal capabilities through a prompt module, a retrieval and domain feedback module, and a conversation module for systematic drug editing. It identifies and manipulates molecular structures for better interpretability in pharmaceutical research. Christofidellis et al. 158 introduced a multi-domain, multi-task language model capable of handling tasks across chemical and natural language domains without specific pretraining for each. Liu et al. 304 developed MolXPT, a unified model that integrates text and molecules for enhanced molecular property prediction and translation, demonstrating robust zero-shot molecular generation capabilities. Zhang et al. 305 introduced PremuNet, a pretrained multi-representation fusion network that enhances molecular property prediction by integrating various molecular data representations. Liu et al. 306 developed GIT-Mol, integrating graph, image, and text data, significantly improving molecular property prediction and molecule generation validity. Gao et al. 307 advanced targeted molecule generation with DockingGA, combining transformer neural networks with genetic algorithms and docking simulations for optimal molecule generation, utilizing Self-referencing Chemical Structure Strings to represent and optimize molecules. Zhou et al. 308 developed TSMMG, a teacher-student LLM for multi-constraint molecular generation, learning from a large set of text-molecule pairs and generating molecules that meet complex property requirements. Zhu et al. 309 proposed 3M-Diffusion, a multi-modal molecular graph generation method that uses natural language descriptions to generate diverse, novel molecular graphs. Gong et al. 310 introduced TGM-DLM, a diffusion model for text-guided molecule generation that overcomes limitations of autoregressive models in generating precise molecules from textual descriptions. Fang et al. 311 developed MolTC, a multimodal framework that effectively integrates graphical information of molecules for improved prediction of molecular interactions. Soares et al. 312 describe MULTIMODAL-MOLFORMER, a multimodal language model that predicts molecular properties by integrating chemical language representations with physicochemical features. It uses a causal multistage feature selection method to identify physicochemical features directly impacting target properties. These features are combined with molecular embeddings from MOLFORMER187, significantly enhancing prediction accuracy for complex tasks like biodegradability and PFAS toxicity.
3.6 Textual scientific LLMs
LLMs are large deep neural network models. Therefore, their superior performance against other smaller models in classical machine learning tasks is somewhat expected. Addressing tasks that seemed impossible to tackle in the past better illustrates LLM’s outstanding capabilities. That is, despite accurately predicting properties given well-structured information such as molecular features and descriptors, LLMs can access less structured data and extract information from simple natural language. For instance, instead of developing a restricted representation, Huang et al. 166 used clinical annotations from the MIMIC-III313 dataset to predict patient readmission. ClinicalBERT implements a pretrained BERT architecture using a combination of masked language modeling and next-sentence prediction training schemes followed by supervised fine-tuning on readmission prediction. Or, instead of pretraining, Zhao et al. 150 fine-tuned LLaMA on epilepsy data using an instruction-following approach to create EpilepsyLLM. Similarly, SciBERT162 and ScholarBERT165 adapted BERT to scientific literature. Beltagy et al. 162 created a tokenizer specific for scientific texts extracted from scientific articles from semantic scholar134 and showed that SciBERT outperforms fine-tuned BERT-base98 in every task investigated. It shows that vocabulary quality plays an important role in the model’s performance. A few years later, Hong et al. 165 pretrained BERT using RoBERTa314 optimizations to improve pretraining performance. ScholarBERT was pretrained on scientific articles from Public.Resource.Org, Inc and further fine-tuned on the tasks used for evaluation. Despite using a much larger dataset, ScholarBERT did not outperform other LLMs trained with narrower domain datasets. Even against SciBERT, another LLM trained on multi-domain articles, ScholarBERT could not perform relatively better in tasks from biomedical, computer science, and material domains. However, ScholarBERT showed significant improvement in the ScienceExamCER315 dataset, which contains semantic class annotations in a training corpus of 3rd to 9th grade science exam questions for named entity recognition (NER) tasks.
On the other hand, instead of pretraining LLMs on a large domain scope, models focused on only one field were also developed. Guo et al. 169 argue that manually curating structured datasets is a sub-optimal, time-consuming, and labor-intensive task. Therefore, they automated data extraction and annotation from scientific papers using ChemDataExtractor316 and their in-house annotation tool317. ChemBERT169 was pretrained using a BERT model to encode chemical reaction information, followed by fine-tuning a NER head. ChemBERT outperformed other models such as BERT98 and BioBERT171 in the product extraction task, presenting an improvement of $\sim$ 6% in precision. For the product role labeling task – given the extracted compound, which role does it play in the reaction? –, it increases precision in $\sim$ 5% when compared to the same models. The increase in performance might be due to training in a narrower domain, which allows the model to learn specific traits from the data instead of being more generalist. MatSciBERT160 and MaterialsBERT167 followed a similar strategy. Gupta et al. 160 developed MatSciBERT by fine-tuning SciBERT162 on the Material Science Corpus (MSC), their curated dataset focused on materials extracted from Elsevier’s scientific papers. MatSciBERT shows an accuracy improvement of $\sim$ 3% compared to SciBERT on classifying an article subject based on its abstract. Differently, MaterialsBERT167 fine-tuned PubMedBERT161 on 2.4M abstracts extracted from material science articles using a masked training approach. MaterialsBERT was then used to generate a contextual encoded representation of the abstract to train a named entity relationship (NER) model. Reporting an increase in precision on their test dataset of $\sim$ 1%, $\sim$ 2%, and $\sim$ 10% for PubMedBERT, MatBERT, and ChemBERT, respectively.
Considerable effort was also employed to develop such models focused on biology tasks. They follow a similar trend of training an LLM model on a large training corpus like Wikipedia, scientific data extracted from a scientific database, and textbooks, followed by fine-tuning specific downstream tasks. Shin et al. 170 pretrained different Megatron-LM318, another BERT-like LLM, models with different sizes, originating the BioMegatron family of models. Considering models with 345M, 800M, and 1.2B parameters and vocabularies with 30k and 50k tokens, they compared the influence of model size on its performance. These models were pretrained using abstracts from the PubMed dataset and full-text scientific articles from PubMed Central (PMC), similar to BioBERT171. The largest 1.2B training started from a BERT-uncased checkpoint to save time. Surprisingly, the largest model did not perform better than the smaller ones. In fact, the 345M parameters using the 50k tokens vocabulary consistently outperformed other models in NER and relation extraction (RE) tasks. Interestingly, the model size seems relevant for the SQuAD319 dataset, showing that LLMs trained on small, domain-specific datasets may lack generalization abilities. BioBERT171 used data from Wikipedia, textbooks, PubMed abstracts, and PMC full-text corpus to pretrain a BERT architecture. Outperforming the original BERT in all benchmarks tested and even SOTA models in some benchmarks (NCBI disease, 2010 i2b2/VA, BC5CDR, BC4CHEMD, BC2GM, JNLPBA, LINNAEUS, Species-800). Peng et al. 320 developed a multi-task BERT evaluated on the BLUE163 benchmark called BlueBERT. The model was pretrained on PubMed abstracts and MIMIC-III313 and fine-tuned on BLUE tasks, achieving similar results as BioBERT in multiple benchmarks. PubMedBERT161, following the approach adopted in SciBERT, developed a domain-specific vocabulary. They used 14M abstracts extracted from PubChem to pretrain PubMedBERT. In addition, they curated and grouped multiple biomedical datasets to develop BLURB, their comprehensive benchmark for biomedical NLP tasks, such as NER, sentence similarity, document classification, and question-answering. Gu et al. 161 showed that PubMedChem considerably outperformed other LLMs in the BLURB benchmark, mainly in PubMedQA and BioQSA datasets. The second best-evaluated model in these datasets was BioBERT, evidencing the importance of domain-specific knowledge to high-performance LLMs.
The models discussed previously in the section use an encoder-only LLM to encode information from scientific data. The learned encoded representation is then further used to make predictions by feeding it to simpler models. A different approach involves training a decoder model. Using a decoder model aims to address the lack of generation ability of the encoder-only models. By implementing a decoder, the model can address a different group of generative tasks, such as question-answering and document classification, using generated labels instead of following a pre-defined set of possible classes. Darwin153 aims to create domain-specific LLMs for natural science by first fine-tuning LLaMA-7B on FAIR, a general Question-Answering (QA) dataset, and then fine-tuning it on scientific QA datasets. Scientific QA instructions were obtained from SciQ321 and generated by the authors using the Scientific Instruction Generation (SIG) model also developed by the authors. Their SIG model is fine-tuned from Vicuna-7B to convert full-text scientific papers into a set of QA pairs. This approach presented an improved performance of DARWIN in several regression and classification benchmarks. Interestingly, Using LLaMA-7B fine-tuned on FAIR only – with no scientific-specific QA fine-tuning – led to similar results in six out of nine benchmarks with metrics available for both models.
In biology, BioGPT156 pretrained a GPT-2 model architecture using 15M abstracts from PubChem corpus. BioGPT was evaluated on four tasks across five benchmarks: end-to-end relation extraction on BC5CDR, KD-DTI, and DDI, question answering on PubMedQA, document classification on HoC, and text generation on the previously listed benchmarks. fine-tuning BioGPT on downstream tasks (except text generation) showed its superiority against other encoder-only LLMs in the biomedical domain, such as BioBERT and PubMedBERT. Focusing on text generation, the authors qualitatively compared BioGPT and GPT-2 generations. They conclude that BioGPT demonstrates better text generation. However, no metric was used, making the comparison difficult to measure. With a similar idea, Wu et al. 154 first pretrained an LLaMA2 model with the MedC-k dataset, consisting of 4.8M academic papers and 30K textbooks. In the sequence, the model was aligned through instruction tuning using the MedC-I dataset, a set of medical QA problems. PMC-LLaMA154 outperforms LLaMa-2 and ChatGPT in multiple biomedical QA benchmarks despite its size $\sim$ 10 times smaller. Interestingly, the authors show that the performance of PMC-LLaMA on MedQA322, MedMCQA323, and PubMedQA324 benchmarks progressively increases as more knowledge is injected, the model size is increased, and more instructions are included in instruction tuning.
3.7 The use of ChatGPT in Chemistry
With the rise of ChatGPT, we review here how many researchers have wanted to test the capability of such an accessible decoder-only LLM. Castro Nascimento and Pimentel 325 wrote the first notable paper on ChatGPT’s impact on Chemistry. The authors emphasize that LLMs, trained on extensive, uncurated datasets potentially containing errors or secondary sources, may include inaccuracies limiting their ability to predict chemical properties or trends. The paper highlighted that while LLMs could generate seemingly valid responses, they lacked true reasoning or comprehension abilities and would perpetuate existing errors from their training data. However, the authors suggested that these limitations could be addressed in the future. The work serves as a benchmark to qualitatively assess improvements in generative pretrained transformers. For example, five tasks were given to ChatGPT (GPT-3). The accuracy for converting compound names to SMILES representations and vice versa was about 27%, with issues in differentiating alkanes and alkenes, benzene and cyclohexene, or cis and trans isomers. ChatGPT found reasonable octanol-water partition coefficients with a 31% mean relative error, and a 58% hit rate for coordination compounds’ structural information. It had a 100% hit rate for polymer water solubility and a 60% hit rate for molecular point groups. Understandably, the best accuracies were achieved with widely recognized topics. The authors concluded that neither experimental nor computational chemists should fear the development of LLMs, nor task automation. Instead, they encouraged the improvement of dedicated AI tools for specific problems, with their integration into research, as facilitators.
The sole use of ChatGPT for chemistry and any apparent advancement in chemistry seems somewhat limited. Humphry and Fuller 326, Emenike and Emenike 327, and Fergus et al. 328 address the use of ChatGPT in chemical education. There are also a few papers that describe the use of ChatGPT on a specific area of chemical research,329, 330, 331 where fine-tuned GPT models are used to direct the synthesis and functional optimization of Metal Organic Framewoks (MOFs), merging computational modeling with empirical chemistry research. Deb et al. 332 provide a thorough assessment of ChatGPT’s effectiveness in material science through various tasks in computational material science, though the perspective is subjective. ChatGPT assisted in computational materials science by providing crystal space groups for simple compounds, suggesting inputs for simulations, refining analysis with specific rules, and identifying resources. The authors highlight ChatGPT’s potential to write code for enhancing processes and as a tool for those without prior expertise, particularly in catalyst development for CO2 capture. We highlight three points about using ChatGPT alone. First, reliable outputs require detailed and accurate input, as Deb et al. 332 showed, where ChatGPT struggled to mine or predict crystal structures. Second, standardized methods for reproducing and evaluating GPT work are still unclear. Third, this work naturally leads to the inference that additional chemical tools or agents are needed to achieve more complex thinking, as described by Bloom’s Taxonomy.333, 334 Bloom’s taxonomy categorizes educational goals into hierarchical levels: Remembering, Understanding, Applying, Analyzing, Evaluating, and Creating. Remembering involves recalling facts, Understanding means restating a concept, and Applying uses a theory in a new situation. Analyzing breaks down a mechanism into parts to understand connections and differences. Evaluating involves making value judgments, and Creating entails building new structures from diverse elements. Currently, LLMs and autonomous agents are limited in replicating higher-level thinking compared to human understanding. To better quantify their capabilities in this area, we therefore recommend adopting a quality metric based on Bloom’s taxonomy. This educational framework can serve as an effective tool for evaluating the sophistication of LLMs and autonomous agents, especially when tackling complex chemical challenges. Bloom’s taxonomy would thereby provide a structured approach to assess higher-order thinking and reasoning.
3.7.1 Automation
The evolution of artificial intelligence in chemistry has fueled the potential for automating scientific processes. For example, in 2019, Coley et al. 335 developed a flow-based synthesis robot proposing synthetic routes and assembling flow reaction systems, tested on medically relevant molecules, and in 2020, Gromski et al. 336 provided a useful exploration on how chemical robots could outperform humans when executing chemical reactions and analyses. They developed the Chemputer, a programmable batch synthesis robot handling reactions like peptide synthesis and Suzuki coupling. In 2021, Grisoni et al. 337 combined deep learning-based molecular generation with on-chip synthesis and testing. The Automated Chemical Design (ACD) framework by Goldman et al. 338 provides a useful taxonomy for automation and experimental integration levels. Thus, automation promises to enhance productivity through increased efficiency, error reduction, and the ability to handle complex problems, as described in several excellent reviews regarding automation in chemistry,339, 340, 341, 342, 343, 344, 345
This increased productivity may be the only possible approach to exploring the vastness of all chemical space. To fully leverage AI in property prediction, inverse design, and synthesis prediction, it must be integrated with automated synthesis, purification, and testing. This automation should be high-throughput and driven by AI-based autonomous decision-making (sometimes called “lights-out” automation). Janet et al. 340 highlighted challenges in multi-step reactions with intermediate purifications, quantifying uncertainty, and the need for standardized recipe formats. They also stated the limitations of automated decision-making. Organa346 addresses some of these challenges. It can significantly reduce physical workload and improve users’ lab experience by automating diverse common lab routine tasks such as solubility assessment, pH measurement, and recrystallization. Organa interacts with the user through text and audio. The commands are converted into a detailed LLM prompt and used to map the goal to the robot’s instructions. Interestingly, Organa is also capable of reasoning over the instructions, giving feedback about the experiments, and producing a written report with the results.
Other limitations exist, like machines’ restriction to pre-defined instructions, their inability to originate new materials, and the lower likelihood of lucky discoveries. Yet, when dedicated tools can be connected to address each step of an automated chemical design, these limitations can be systematically addressed through advancements in LLMs and autonomous agents, discussed in the next section.
4 LLM-based autonomous agents
The term “agent” originates in philosophy, referring to entities capable of making decisions347. Hence, in artificial intelligence, an “agent” is a system that can perceive its environment, make decisions, and act upon them in response to external stimuli348. Language has enabled humans to decide and act to make progress in response to the environment and its stimuli, and so LLMs are naturally ideal for serving as the core of autonomous agents. Thus, in agreement with Gao et al. 349, we define a “language agent” as a model or program (typically based on LLMs) that receives an observation from its environment and executes an action in this environment. Here, environment means a set of tools and a task. Hence, “LLM-based autonomous agents” refer to language agents whose core is based on an LLM model. Comprehensive analyses of these agents are available in the literature350, 348, 349, but this section highlights key aspects to prepare the reader for future discussions.
There is no agreed definition of the nomenclature to be used to discuss agents. For instance, Gao et al. 349 created a classification scheme that aims to group agents by their autonomy in biological research. This means a level 0 agent has no autonomy and can only be used as a tool, while a level 3 agent can independently create hypotheses, design experiments, and reason. Another previously suggested framework highlights agents’ construction, application, and evaluation as a parallel between agents and neural networks. Following this perspective, Wang et al. 350 categorizes agent components into four modules: profiling, memory, planning, and action. In contrast, Weng 351 also identifies four elements — memory, planning, action, and tools — but with a different emphasis. Meanwhile, Xi et al. 348 proposes a division into three components: brain, perception, and action, integrating profiling, memory, and planning within the brain component, where the brain is typically an LLM. Recently, Sumers et al. 352 proposed Cognitive Architectures for Language Agents (CoALA), a conceptual framework to generalize and ease the design of general-purpose cognitive language agents. In their framework, a larger cognitive architecture composed of modules and processes is defined. CoALA defines a memory, decision-making, and core processing module, in addition to an action space composed of both internal and external tools. While internal tools mainly interact with the memory to support decision-making, external tools compose the environment, as illustrated in Figure 5. Given a task that initiates the environment, the “decision process” runs continuously in a loop, receiving observations and executing actions until the task is completed. For more details, read Reference [352].
Figure 5: Agent’s architecture as defined in this review. According to our definition, an agent is composed of a central program (typically an LLM and the code to implement the agent’s dynamic behavior) and the agent modules. The agent continuously receives observations from the environment and decides which action should be executed to complete the task given to it. Here, we define the agent as the set of elements whose decision is trainable, that is, the LLM, the agent code, the decision process, and the agent modules. Given a task, the agent uses the agent modules (memory, reasoning, planning, profiling) and the LLM to decide which action should be executed. This action is executed by calling a tool from the environment. After the action is executed, an observation is produced and fed back to the agent. The agent can use perception to receive inputs in different modalities from the environment. A) Description of agent modules, B) illustration of the agent architecture, C) illustration of the environment components, D) description of tools elements present in the environment.
In this review, we define autonomous agent systems as models (typically LLMs) that continuously receive observations and execute actions within an environment, as described by Gao et al. 349. Nevertheless, in contrast to CoALA349, we will rename “internal tools” as “agent modules” and “external tools” simply as “tools”, for clarity. The agent consists of trainable decision-making components such as the LLM itself, policy, memory, and reasoning scheme. In contrast, the environment comprises non-trainable elements like the task to be completed, Application Programming Interface (API) access, interfaces with self-driving labs, dataset access, and execution of external code. By referring to decision-making components as agent modules, we emphasize their inclusion as parts of the agent. By referring to non-trainable elements as tools, we highlight their role as part of the environment. We discuss five main types of actions. As shown in Figure 5, three of the five, memory, planning and reasoning, and profiling are agent modules. The remaining two actions, tools and perception, are part of the environment. Since the perception is how the agent interacts with the environment and is not a trainable decision, we therefore included it as part of the environment.
4.1 Memory module
The role of the memory module is to store and recall information from past interactions and experiences to inform future decisions and actions. There are multiple types of memory in agents, namely sensory memory, short-term memory, and long-term memory. A major challenge in using agents is the limited context window, which restricts the amount of in-context information and can lead to information loss, thereby impacting the effectiveness of short-term and long-term memory. Solutions involve summarizing memory content353, compressing memories into vectors354, 355, 356, and utilizing vector databases357 or combinations thereof358, with various databases available such as ChromaDB, FAISS, Pinecone, Weaviate, Annoy, and ScaNN.359 Addressing these challenges to enhance agent memory continues to be a significant area of research360. Sensory, or procedural, memory is knowledge embedded into the model’s parameters during pretraining and/or in heuristics implemented into the agent’s code. Short-term, or working, memory includes the agent’s finite knowledge during a task, incorporating interaction history and techniques like in-context learning361 (ICL), which leverages the limited input’s context length for information retention. Long-term memory involves storing information externally, typically through an embedded vector representation in an external database. In the original CoALA349 paper, long-term memory is further divided into two different types of long-term memory: episodic, which registers previous experiences, and semantic, which stores general information about the world.
4.2 Planning and reasoning modules
The planning and reasoning module is made of two components. Planning involves identifying a sequence of actions required to achieve a specified goal. In the context of language agents, this means generating steps or strategies that the model can follow to solve a problem or answer a question, which can be enhanced with retrieval from previous experiences,362 and from feedback from post-execution reasoning363, 364. Reasoning refers to the process of drawing conclusions or making decisions based on available information and logical steps. For example, there are studies that demonstrate the benefits of LLM reasoning for question answering, where new context tokens can be integrated in a step-by-step way to guide the model towards more accurate answers365, 366, 367, 368, 369, 370. One popular reasoning strategy is Chain-of-Thought (CoT),367, 371, 372, 91, 373, 374 a reasoning strategy which substantially boosts QA performance by generating intermediate reasoning steps. CoT thereby reduces hallucinations and enhances interpretability, as demonstrated by improved results in models like PaLM375 and GPT-3 with benchmarks like GSM8K376, SVAMPs377, and MAWPS378.
In advanced reasoning, final tasks are often decomposed into intermediary ones using a cascading approach, similar to Zero-shot-CoT366 and RePrompt363. However, while CoT is considered as single-path reasoning, CoT extensions like Tree-of-Thoughts369, Graph-of-Thoughts379, Self-consistent CoT368, and Algorithm-of-Thoughts380 offer multi-path reasoning. Furthermore, other models have pitted multiple agents against each other to debate or discuss various reasoning paths,381, 382, 383 while others use external planners to create plans384, 385. A feedback step during the execution of the plan was a further extension of the CoT ideas; this enables agents to refine their actions based on environmental responses adaptively, which is crucial for complex tasks.386, 387
Another interesting reasoning scheme is the Chain-of-Verification(CoVe), 364 where once an answer is generated, another LLM is prompted to generate a set of verification questions to check for agreement between the original answer and the answers to the verification questions such that the final answer can be refined. The ReAct369 – Reason+Act – model proposes adding an observation step after acting. This means the LLM first reasons about the task and determines the necessary step for its execution, it performs the action and then observes the action’s result. Reasoning on that result, it can subsequently perform the following step. Similarly, Reflexion91 also implements a reasoning step after executing an action. However, Reflexion implements an evaluator and self-reflection LLMs to not only reason about each step but also to evaluate the current trajectory the agent is following using a long-term memory module. As the context increases, it may become challenging for agents to deal with the long prompt. Aiming to solve this issue, the Chain-of-Agents (CoA)388 extends reasoning schemes that leverage multi-agent collaboration to reason over long contexts. This framework employs workers and manager agents to process and synthesize information to generate the final response. CoA demonstrated improvements of up to 10% when compared against an RAG baseline.
ReAct and Reflexion are closed-ended approaches where the agent starts with all the tools and must determine which to use. To address more open-world challenges, Wang et al. 389 introduced the Describe, Explain, Plan, and Select (DEPS) method, which extends this approach. Lastly, human inputs can also be used to provide feedback to the agent. Providing feedback using a human-in-the-loop approach is particularly interesting in fields where safety is a main concern.
4.3 Profiling module
LLMs can be configured to perform in specific roles, such as coders, professors, students, and domain experts, through a process known as profiling. Language agents can thus incorporate the profile through the LLM or through the agent code. The profiling approach involves inputting psychological characteristics to the agent, significantly impacting its decision-making process 390, 391, 392, 393. Profiling enables the creation of multi-agent systems that simulate societal interactions, with each agent embodying a unique persona within the group 362, 394. The most prevalent technique for profiling, called “handcrafting”, requires manually defining the agent’s profile, often through prompts or system messages 395, 396. While profiling can also be automated with LLMs 397, that automation method may only be suited for generating large numbers of agents since it offers less control over their overall behavior. An interesting application of profiling is the development of agent sets that reflect demographic distributions 398.
4.4 Perception
Perception is an analog to the human sensory system, which interprets multimodal information such as text, images, or auditory data, transforming it into a format comprehensible by LLMs, as demonstrated by SAM399, GPT4-V400, LLaVa401, Fuyu8B109, and BuboGPT402. In our proposed architecture, the perception is responsible for converting the task and the observations to a data representation that can be understood by the agent. Moreover, advancements in LLMs have led to the development of even more versatile models, such as the any-to-any Next-GPT110 and the any-to-text Macaw-LLM403. Employing multimodal LLMs, which are introduced above, in decision-making processes can simplify perception tasks for agents, with several studies exploring their use in autonomous systems404, 405.
4.5 Tools
In our proposed definition (see Figure 5b), tools or actions are part of the environment. The agent can interact with this environment by deciding which action to execute through the decision-making process. The set of all possible actions that can be selected is also known as the “action space”.
The decision process is composed of three main steps: proposal, evaluation, and selection. During the proposal, one or more action candidates are selected using reasoning369, code structures406, 362, or simply by selecting every tool available407.365, 408, 368 The evaluation process consists of evaluating each selected action according to some metric to predict which action would bring more value to the agent. Lastly, the action is selected and executed.
Given that pretrained parameters (sensory memory) are limited, the model must use tools for complex tasks in order to provide reliable answers. However, LLMs need to learn how to interact with the action space and how and when to use those tools most accurately.409 LLMs can be pretrained or fine-tuned with examples of tool use, enabling them to operate tools and directly retrieve tool calls from sensory memory during a zero-shot generation.410 Recent studies investigate this approach, particularly focusing on open-source LLMs.411, 412, 413
As foundational AI models become more advanced, their abilities can be expanded. It was shown that general-purpose foundation models can reason and select tools even with no fine-tuning. For example, MRKL414 (pronounced “miracle”) implements an extendable set of specialized tools known as neuro-symbolic modules and a smart “router” system to retrieve the best module based on the textual input. These neuro-symbolic modules are designed to handle specific tasks or types of information, equipped with built-in capabilities and task-relevant knowledge. This pre-specialization allows the model to perform domain-specific tasks without needing a separate, domain-specific dataset. This design addresses the problem of LLMs lacking domain-specific knowledge and eliminates the need for the costly and time-consuming LLM fine-tuning step, using specialized data annotation.415 The router can receive support from a reasoning strategy to help select the tools415 or follow a previously created plan389. Recent advances have shown that LLMs can develop new tools of their own416, 417, 418, enabling agents to operate, as needed, in dynamic and unpredictable “open-worlds”, on unseen problems as illustrated by Voyager406. This capability allows agents to evolve and improve continually.
5 LLM-Based autonomous agents in scientific research
The previous section introduced key concepts relevant to any description of the development of autonomous agents. Here, we now focus on which agents were developed for scientific purposes, and ultimately for chemistry. Previous sections of this review have discussed how LLMs could be powerful in addressing challenges in molecular property prediction, inverse design, and synthesis prediction. When we consider the value of agents in chemistry and the ability to combine tools that, for example, search the internet for established synthetic procedures, look up experimental properties, and control robotic synthesis and characterization systems, we can see how autonomous agents powerfully align with the broader theme of automation, which will lead to an acceleration of chemical research and application.
Table 4: Scientific LLM systems and agents. We identify the studies we classified as an agent with the icon [Uncaptioned image]419 and multi-agent systems with the icon [Uncaptioned image][Uncaptioned image]. [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image]420 mean the agent bases his behavior on sensory, short, and long memory components, respectively. Besides the textual capabilities of LLM-based agents, [Uncaptioned image]421 and [Uncaptioned image]422 mean the agent has additional audio and visual perception, respectively.
Agent Memory Planning Reasoning Action Release
LLaMP423[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Tools for database access, literature search, and atomistic simulations 2024.06
CRISPR-GPT424[Uncaptioned image][Uncaptioned image] [Uncaptioned image] Tool for gene editing experiments design 2024.04
TAIS425[Uncaptioned image][Uncaptioned image] [Uncaptioned image] Tools for gene expression data analysis 2024.02
SciAgent426[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Trained Mistral for tool usage. Evaluated it using MathToolBench’s tools 2024.02
STORM427 [Uncaptioned image][Uncaptioned image] Article writing using retrieval from multi-LLM conversations and pre-generated outline 2024.02
Völker et al. 428 [Uncaptioned image] Regression with ICL and text retrieval 2024.02
ProtAgent429 [Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image] Tools for proteins information retrieval, analyzing, de novo design, and 3D folded structure generation 2024.01
PaperQA430[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Tools to search the scientific literature, gather evidence, and answer questions 2023.12
WikiCrow431[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Uses PaperQA as a tool 2023.12
Coscientist46[Uncaptioned image] [Uncaptioned image] Tools for running Python code, web-searching, and interacting with lab equipment 2023.12
Eunomia432[Uncaptioned image] [Uncaptioned image] Tools for literature and dataset searching and a chain-of–verification loop 2023.12
CALMS433[Uncaptioned image] [Uncaptioned image] Tools for using the Materials Project API, designing experiments, and using a hardware API to perform the experiment 2023.12
CoQuest434[Uncaptioned image] [Uncaptioned image] Research question generations and tools for literature visualization using a graph organization 2023.10
eXpertAI286[Uncaptioned image] [Uncaptioned image] Tools for applying XAI methods 2023.11
BioPlanner435[Uncaptioned image] [Uncaptioned image] Tool for protocol searching in the BioProt dataset 2023.10
IBM ChatChem436[Uncaptioned image] [Uncaptioned image] Tools for cheminformatics and accessing GT4SD and HuggingFace models 2023.09
ChatMOF437[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Tools for database search, property prediction, and MOF’s structure generation 2023.08
AmadeusGPT45[Uncaptioned image][Uncaptioned image] [Uncaptioned image][Uncaptioned image] Tools for writing and executing code for computer vision, machine learning, and spatial-temporal reasoning 2023.07
i-Digest438 [Uncaptioned image] [Uncaptioned image] Uses the whisper model to process audio transcription from classes and write summaries and following up questions 2023.06
BOLLaMa 439 [Uncaptioned image] Implements an LLM interface to ease the usage of their BO code 2023.06
text2concrete440 [Uncaptioned image] Uses ICL to predict compressive strength from concrete formulation 2023.06
MAPI_LLM441[Uncaptioned image] [Uncaptioned image][Uncaptioned image] Database access and LLM prediction using ICL 2023.06
BO-LIFT442 [Uncaptioned image] Regression using ICL and text retrieval 2023.04
ChemCrow45[Uncaptioned image] [Uncaptioned image] Molecular, cheminformatics, search and critique tools 2023.04
It was Hocky and White 443 who discussed the early stages of models that could automate programming and, hence, the expected impacts in chemistry. Then, early work by White et al. 444 applied LLMs that could generate code to a benchmark set of chemical problems. In that case, not only were LLMs demonstrated to possess a notable understanding of chemistry, based on accurate question answering, but White et al. 444 imagined a potential to use them as base models to control knowledge augmentation and a variety of other tools. Thus, these LLMs could be used to execute routine tasks, optimize procedures, and enhance the retrieval of information from scientific literature across a range of scientific domains. To the best of our knowledge, this is the first review of autonomous agents in chemistry that have evolved since these two visionary conceptual perspectives. A deeper exploration follows below. One driving motivation for the need to augment LLMs with a more pertinent and dedicated knowledge base is the need to circumvent problems of a limited context prompt window, and the restriction that once an LLM is trained, any new information is beyond it’s reach since it necessarily has fallen outside its corpus of training data. Furthermore, LLMs are also known to hallucinate. Their predictions are probabilistic and, in science, if experimental evidence is available, then there is great value in building from known domain-specific information. Some improved prompt engineering can aid in the generation of results that are more likely to be accurate, but the use of autonomous agents may solve such problems completely in this next phase of AI in chemistry. In fact, even adding one or two components when building an agent, as opposed to a whole suite, has shown some significant gains.
Building on this foundation, Ramos et al. 442 illustrated that LLMs could directly predict experimental outcomes from natural language descriptions, a technique they incorporated into a Bayesian optimization (BO) algorithm to streamline chemical processes. Because ICL was used, their approach did not require additional model training or finetuning, thereby greatly simplifying the optimization algorithm and removing the need to re-train or update models as data is gathered. Recently, Kristiadi et al. 445 demonstrated that similar results could also be achieved with a much smaller domain-specific model while performing parameter-efficient fine-tuning (PEFT) instead of ICL. Ranković and Schwaller 446 also explored BO using natural language. They used an LLM to encode chemical reaction procedures, described using natural language, and then trained a Gaussian process (GP) head to predict the reaction yield from the latent encoded representation of the procedure. This effective transfer-learning approach minimized training time by keeping the LLM frozen while updating only the MLP. Extending these ideas, Völker et al. 428 suggested sampling multiple model completions and adding a verifier model to select the next best procedure to continue the BO algorithm. They also used ICL and a short-term memory component to optimize alkali-activated concrete mix design. The authors thus underscored how augmenting with knowledge-driven design can outperform traditional data-driven design. These are examples of how agent-based approaches can execute complex optimization algorithms in a step-by-step approach, directly contributing to automation and leading to faster and more efficient experimental design.
To better promote new ideas regarding AI in scientific research, Jablonka et al. 447 organized a hackathon in March 2023. During this one-day hackathon, 14 innovative projects were developed to address chemical problems centered around predictive modeling, automation, knowledge extraction, and education. Below, we highlight agent-based approaches from that Hackathon. First, MAPI_LLM441 is an agent with access to the Materials Project API (MAPI) database that receives a query asking for a property of a material and then retrieves the relevant information from the dataset. If the material is not available on MAPI, the agent can search for similar materials and use in-context learning (ICL) to provide a prediction of the requested property. Additionally, MAPI_LLM also has a reaction module for synthesis proposal. Second, Rankovic et al. 439 used LLMs to make BO algorithms more accessible to a broader group of scientists; BOLLaMa implements a natural language interface to easily interact with BO software developed by their group448. Third, and similar to Ramos et al. 442 and Ranković and Schwaller 446 who employed LLMs in BO, Weiser et al. 449 focused on genetic algorithms (GA), a different optimization algorithm. In GA, pieces of information are stochastically combined and evaluated to guide the algorithm during the optimization. For chemistry, these pieces are often molecular fragments that are combined to compose a final whole molecular structure. Thus, Weiser et al. 449 used LLMs to implement common GA operators under the hypothesis that LLMs can generate new combined molecules better than random cross-over due to their sensory memory. Fourth, InsightGraph450 can draw general relationships between materials and their properties from JSON files. Circi and Badhwar 450 showed that LLMs can understand the structured data from a JSON format and reorganize the information in a knowledge graph. Further refinement of this tool could automate the process of describing relationships between materials across various scientific reports, a task that remains labor-intensive today. Fifth, Kruschwitz et al. 440 used ICL and LLMs to accurately predict the compressive strength of concrete formulations; Text2Concrete achieved predictive accuracy comparable with a Gaussian process regression (GPR) model, with the advantage that design principles can be easily added as context. This model was successfully applied in a BO algorithm following Ramos et al. 442 approach.428 For education purposes, multiple authors have raised the discussion about how LLMs can be used to support educators’ and instructors’ daily work.451, 452, 453, 454, 455 Finally, in this direction, Mouriño et al. 438 developed i-Digest, an agent whose perception module can understand audio tracks and video recordings. These audio recordings are transcribed to text using the Whisper456 model, and therefore, i-Digest is a digital tutor that generates questions to help students test their knowledge about the course material. These are just a few examples to showcase the capabilities of AI systems to innovate and generate solutions rapidly.
More recently, Ma et al. 426 showed that agents can be trained to use tools. SciAgent426 was developed under the premise that finetuning LLMs for domain-specific applications is often impractical. Nevertheless, the agent can be fine-tuned with a set of tools that will enable them to perform well in a domain-specific task. These tools, typically Python functions, enable SciAgent to plan, retrieve, and use these tools to facilitate reasoning and answer domain-related questions effectively. The benchmark developed for SciAgent, known as SciToolBench, includes five distinct domains, each equipped with a set of questions and corresponding tools. The development of its retrieval and planning modules involved finetuning different LLMs on the MathFunc benchmark, resulting in a notable performance improvement of approximately $\sim$ 20% across all domains in SciToolBench compared to other LLMs.
5.1 Agents for literature review
Another fantastic opportunity for automation in the sciences is associated with high-quality literature review, a pivotal aspect of scientific research that requires reading and selecting relevant information from large numbers of papers, and thereby distilling the current state of knowledge relevant to a particular research direction. This extremely time-consuming task is being revolutionized by advanced AI tools designed to automate and enhance such analysis and summarization.
PaperQA introduces a robust model that significantly reduces misinformation while improving the efficiency of information retrieval. This agent retrieves papers from online scientific databases, reasons about their content, and performs question-answering (QA) tasks. Its mechanism involves three primary components—“search”, “gather_evidence”, and “answer_question” and the authors adapted the Retrieval-Augmented Generation (RAG)457 algorithm to include inner loops on each step. For instance, PaperQA can perform multiple rounds of search and gather_evidence if, with reflection, not have enough evidence has been acquired to successfully answer_question. Further validating its capabilities, a new benchmark called LitQA demonstrates that PaperQA not only meets but exceeds human performance in solving complex, real-world scientific questions. By applying the RAG technique to full-text scientific papers, PaperQA sets a new standard in QA capabilities, achieving human-like performance in curated datasets without hallucination or selecting irrelevant citations.430 To further investigate PaperQA performance, the authors developed a new benchmark called LitQA, which designs tasks that mimic the complexity of real scientific inquiry. The benchmark comprises 50 multiple-choice questions sourced from biomedical papers published post-September 2021, ensuring they were not part of the LLM training data. In this setting, PaperQA demonstrated a precision rate of 87.9% and surpassed typical human accuracy, scoring 69.5% compared to 66.8%.430
Building on top of PaperQA, WikiCrow exemplifies the practical application of AI in generating concise and relevant Wikipedia-style summaries. The authors show that while 16% of a human-created Wikipedia article comprises irrelevant statements, WikiCrow displays irrelevant information only 3% of the time. Their system also added 5% more correct citations when compared with original articles. Moreover, thanks to its foundation in the PaperQA framework430, WikiCrow achieves remarkable cost-efficiency. The authors estimate that WikiCrow can accomplish in a few days what would take humans approximately 60,000 hours, or about 6.8 years, thereby underscoring its ability to rapidly produce extensive scientific content. This efficiency exemplifies the reliability and transformative potential of AI in content creation 431.
Following a different approach, the STORM model also addressed the problem of writing Wikipedia-like summaries, where the STORM acronym represents the Synthesis of Topic Outlines through Retrieval and Multi-perspective questions.427 This approach implements a two-step procedure. First, STORM retrieves multiple articles on a topic and uses an LLM to integrate various perspectives into a cohesive outline. Second, this outline is used to write each section of the Wikipedia-like summary individually. To create the outline, multiple articles discussing the topic of interest are retrieved by an “expert” LLM, which processes each one to create $N$ perspectives. Each perspective is then fed to a “writer” LLM, and a conversation is initiated between writer and expert. Finally, the $N$ conversations are used to design the final outline. The outline and the set of references, accessed by RAG, are given to the writer LLM. The writer LLM is prompted to use these inputs to generate each section of the article sequentially. Following this, all sections are merged and refined to eliminate redundancies and enhance coherence. Upon human evaluation, STORM is reported to be $\sim$ 25% more organized and present $\sim$ 10% better coverage when compared to a pure RAG approach. However, it was also less informative than human-written Wikipedia pages, and STORM presented a transfer of internet-borne biases, producing emotional articles, which is a major concern.
Ultimately, the idea of an “AI scientist” who can generate new, relevant research questions (RQ) has been pursued. Aligned with this vision, Liu et al. 434 developed CoQuest, partially automating the brainstorming for new RQs process. This system uses a human-computing interface (HCI) to allow the agent to create new RQs that can be further enhanced by human feedback. They developed two strategies for RQ generation: breadth-first, where the agent generates multiple RQs simultaneously following the original user’s prompt, and depth-first, where multiple RQs are created sequentially, building on the top of the previously generated RQ. For each RQ generation, the agent implements a ReAct369 framework with tools for literature discovery, hypothesis proposition, refinement, and evaluation. Upon evaluation of 20 HCI doctoral researchers by a post-interaction survey, the breadth-first approach was preferred by 60% of the evaluators. Interestingly, despite the evaluators’ report that the breadth-first approach gave them more control and resulted in more trustworthy RQs, the depth-first had better scores for novelty and surprise. This difference might be caused by the fact that the depth-first uses its own RQ to iterate. This process can introduce new keywords not considered by the users.
5.2 Agents for chemical innovation
Transitioning from literature synthesis to practical chemistry applications, we next explore how LLM-based agents have proven their capabilities to revolutionize routine chemical tasks towards an acceleration of molecular discovery and scientific research. Agents are flexible entities capable of developing prompt-specific workflows and executing a plan toward accomplishing a specific task. ChemCrow45 introduced a significant shift in how LLMs would be applied in chemistry, given that LLMs alone do not access information outside of their training data nor can they directly perform chemistry-related tasks.
By augmenting LLMs with common chemical tools, computational or robotic, ChemCrow automates a broad spectrum of routine chemical tasks, demonstrating a significant leap in LLM applicability. Under human evaluation, ChemCrow consistently outperformed GPT-4, achieving an accuracy score of 9.24/10 compared to 4.79/10.45 The developers of ChemCrow have also considered the ethical implications and potential risks associated with its capabilities. ChemCrow’s high potential could be misused and exploited for malicious objectives, and therefore the authors have implemented safety checks and guidelines to prevent such misuse, or “dual usage”. Additionally, they acknowledge that ChemCrow, relying on an LLM, may not always provide completely accurate answers due to gaps in its chemical knowledge. As such, they recommend careful and responsible use of the tool, along with thorough scrutiny of its outputs. In summary, while ChemCrow presents a powerful new chemical assistant45, oversight of its use is required, and this agent’s access to tools has been deliberately limited to enhance security and avoid misuse.
Similarly to ChemCrow45, Chemist-X458 uses RAG to get up-to-date literature information and use it to reliably solve user’s questions. Nevertheless, Chemist-X focuses on designing chemical reactions to achieve a given molecule. It works in three phases: (1) First, the agent searches molecule databases for similar molecules, then (2) there is a look-up for online literature searching for chemical reactions capable of converting the list of similar molecules in the target. Lastly, (3) machine learning models are used to propose the reaction conditions. To validate their agent, the authors used Chemist-X to design an HTS experiment aiming to produce 6-(1-methyl-1H-indazol-4-yl), resulting in a maximum yield of 98.6%.
On the other hand, the Coscientist46 system exemplifies the integration of semi-autonomous robots in planning, conceiving, and performing chemical reactions with minimal human intervention. At its core, the system features a main module named ‘PLANNER’, which is supported by four submodules. These submodules, or tools, are responsible for performing actions such as searching the web for organic synthesis, executing Python code, searching the hardware documentation, and performing a reaction in an automated lab.46 Utilizing this framework, the Coscientist successfully conducted two types of chemical coupling reactions, Suzuki-Miyaura and Sonogashira, in a semi-automated fashion, with manual handling of initial reagents and solvents. Additionally, Coscientist was also used to optimize reaction conditions. In contrast to Ramos et al. 442, who used LLMs within a Bayesian Optimization (BO) algorithm as a surrogate model, Boiko et al. 46 approached the optimization task as a strategic “game” aimed at maximizing reaction yield by selecting optimal reaction conditions. This demonstrates the ability of GPT-4 to effectively reason about popular chemical reactions – possibly via comprehensive coverage in pretraining. The authors have indicated that the code for their agent will be released following changes in U.S. regulations on AI and its scientific applications. At the time of writing, the code remains unreleased, but a simple example that calculates the square roots of random numbers has been provided to illustrate their approach.46. These examples underscore the transformative role of LLMs in enhancing and automating chemical processes, which will likely accelerate chemical discovery.
Automated workflows in protein research have also been explored. ProtAgent429 is a multi-agent system designed to automate and optimize protein design with minimal human intervention. This system comprises three primary agents: Planner, Assistant, and Critic. The Planner is tasked with devising a strategy to address the given problem, the Assistant executes the plan using specialized tools and API calls, and the Critic supervises the entire process, providing feedback and analyzing outcomes. These agents collaborate through a dynamic group chat managed by a fourth agent, the Chat Manager. Tasks executed by this team include protein retrieval and analysis, de novo protein design, and conditioned protein design using Chroma459 and OmegaFold460.
Similarly to ProtAgent, Liu et al. 425 created a team of AI-made scientists (TAIS) to scientific discovery without human intervention. However, their agents have roles analog to human roles in scientific discovery, for instance, project manager, data engineer, code reviewer, statistician, and domain expert. While in ProtAgent429 agents interact through the Chat Manager only, TAIS425 enables AI scientists to interact between themselves directly using pre-defined collaboration pipelines. To evaluate TAIS, the authors curated the Genetic Question Exploration (GenQEX) benchmark, which consists of 457 selected genetic data questions. As a case study, the authors show TAIS’s answer to the prompt ‘‘What genes are associated with Pancreatic Cancer when considering conditions related to Vitamin D Levels?’’. The system identified 20+ genes with a prediction accuracy of 80%.
Innovation can also be achieved by looking into data from a different point-of-view to get new insights. Automating querying databases was investigated by Ramos et al. 441 with a ReAct agent with access to the MAPI dataset. This concept was extended by Chiang et al. 423. LLaMP423 is a RAG-based ReAct agent that can interact with MAPI, arXiv, Wikipedia, and has access to atomistic simulation tools. The authors showed that grounding the responses on high-fidelity information (a well-known dataset) enabled the agent to perform inferences without fine-tuning.
The agents in chemistry, as exemplified by ChemCrow45 and Coscientist46, highlight a significant shift towards automation and enhanced efficiency in molecular discovery and scientific research. These systems demonstrate the potential of integrating LLMs with chemical tools and automation frameworks, achieving impressive accuracy and effectiveness in tasks ranging from routine chemical operations to complex reaction optimizations. Similarly, ProtAgent429 and TAIS425 systems showcase the versatility of multi-agent frameworks in automating protein design and genetic research, pushing the boundaries of what AI-driven scientific discovery can achieve. These studies collectively showcase the incredible potential of agents in chemical and biological research, promising automation of routine tasks, easing the application of advanced techniques and analyses, and accelerating discoveries. However, they also underscore the necessity for meticulous oversight and responsible development to harness their full potential while mitigating risks.461
5.3 Agents for experiments planning
Building on the capabilities of ChemCrow and Coscientist in automating chemistry-related tasks, recent advances have focused on bridging the gap between virtual agents and physical laboratory environments. For example, Context-Aware Language Models for Science (CALMS)433, BioPlanner435, and CRISPR-GPT424 focus on giving support to researchers with wet-lab experimental design and data analysis.
CALMS433 focuses on improving laboratory efficiency through the operation of instruments and management of complex experiments, employing conversational LLMs to interact with scientists during experiments. In addition, this agent can perform actions using lab equipment after lab equipment APIs have been provided to the agent as tools. CALMS was designed to enhance instrument usability and speed up scientific discovery, providing on-the-spot assistance for complex experimental setups, such as tomography scans, and enabling fully automated experiments. For instance, its capability was showcased through the operation of a real-world diffractometer. Although CALMS excelled in several tasks, a comparison between GPT-3.5 and Vicuna 1.5 revealed Vicuna’s limitations in handling tools.
In contrast, BioPlanner435 significantly improves the efficiency of scientific experimentation by creating pseudocode representations of experimental procedures, showcasing AI’s capacity to streamline scientific workflows. Therefore, Rather than interacting directly with lab equipment through APIs, BioPlanner creates innovative experimental protocols that can be expanded upon within a laboratory setting. The initial step in BioPlanner’s process involves assessing the capability of LLMs to produce structured pseudocode based on detailed natural language descriptions of experimental procedures. In testing, BioPlanner successfully generated correct pseudocode for 59 out of 100 procedures using GPT-4, although the most common errors involved omitted units. Afterward, the authors used BioPlanned to generate a procedure for culturing an E.coli bacteria colony and storing it with cryopreservation, which ran successfully.
Focusing on gene editing experiments, CRISPR-GPT424 is an agent developed to design experiments iteratively with constant human feedback. CRISPR-GPT424 aims to bridge the gap for non-experts by simplifying this process into manageable steps solvable by an LLM with access to useful tools. This agent operates in three modes based on user prompts: “Meta mode” provides predefined pipelines for common gene-editing scenarios; “Auto mode” uses the LLM to plan a sequence of tasks; and “Q&A mode” answers general questions about the experimental design. The authors demonstrate that based on human evaluations, CRISPR-GPT outperforms GPT-3.5 and GPT-4 in accuracy, reasoning, completeness, and conciseness. Additionally, they applied CRISPR-GPT to design real-world experiments for knocking out TGFBR1, SNAI1, BAX, and BCL2L1 in the human A375 cell line, achieving an editing efficiency of approximately 70% for each gene.
Following the ideas of developing agents for automating experimental protocol generation, Ruan et al. 462 created a multi-agents system composed of 6 agents: Literature Scouter, Experiment Designer, Hardware Executor, Spectrum Analyzer, Separation Instructor, and Result Interpreter. The Large Language Models-based Reaction Development Framework (LLM-RDF)462 automates every step of the synthesis workflow. While other studies focus on the literature review463, 464, 465, HTS458, and reaction optimization466, 46, LLM-RDF can support researchers from literature search until the product purification. Using this system, the authors showed they could design a copper/TEMPO catalyzed alcohol oxidation reaction, optimize reaction conditions, engineer a scale-up, and purify the products, obtaining a yield of 86% and a purity >98% while producing 1 gram of product.
Interestingly, despite acting on different fields and presenting different goals (while CALMS433 and LLM-RDF462 can autonomously use laboratory equipment, BioPlanner435 and CRISPR-GPT424 propose a protocol that needs to be carried out by a human), all studies implemented a “human-in-the-loop” approach. Maintaining the researcher as a vital piece of the development is a way to enhance reliability and mitigate agent limitations, such as errors and hallucinations. This approach also addresses risk and dual usage problems once a human can decide whether the agent’s suggestions are safe. 350, 467 On a slightly different direction, Organa346 fully automates the laboratory workload while providing feedback to the researcher and producing reports with the results, as discussed on Section 3.7.1.
5.4 Agents for automating cheminformatics tasks
Cheminformatics consists of applying information technology techniques to convert physicochemical information into knowledge. The process of solving cheminformatics problems commonly involves retrieving, processing, and analyzing chemical data.468 Getting inspiration from ChemCrow469 ideas, Chemistry Agent Connecting Tool Usage to Science (CACTUS)470 focused on assisting scientists by automating cheminformatics tasks. CACTUS automates the applications of multiple cheminformatics tools, such as property prediction and calculation, while maintaining the human-in-the-loop for molecular discovery. The authors investigated the performance of a diverse set of open-source LLMs, where Gemma-7B and Mistral-7B demonstrated superior performance against LLaMA-7B and Falcon-7B. In addition, the authors reported that adding domain-specific information in the prompt to align the agent to chemistry problems considerably increases model’s performance. For instance, predicting drug-likeness with a Gemma-7B agent improves the accuracy of $\sim$ 60% when aligning the agent. Prompt alignment improved the prediction of all properties they studied.
Further illustrating the versatility of AI in scientific research and domain-specific tools usage is ChatMOF,437 which focuses on the prediction and generation of Metal-Organic Frameworks (MOFs). ChatMOF integrates MOF databases with its MOFTransformer471 predictor module, thereby showcasing the innovative use of genetic algorithms in guiding generative tasks from associated predictions. The authors showed that ChatMOF achieved an accuracy of $\sim$ 90% in search and prediction tasks while generative tasks have an accuracy of $\sim$ 70%. The genetic algorithm used by ChatMOF allows for the generation of a diverse array of MOF structures, which can be further refined based on specific properties requested by users. For instance, when prompted to, “generate structures with the largest surface area”, the system initially generated a broad distribution of structures with surface area centered in 3784 m2/g, and the GA evolves it to a narrower distribution with a peak at 5554 m2/g after only three generations. It is important to note that even though ChatMOF has access to a dataset of experimental values for MOFs, language model predictions guide their GA, and no further validation has been made. Lastly, Ansari and Moosavi 432 developed Eunomia, another domain-specific autonomous AI agent that leverages existing knowledge to answer questions about materials. Eunomia432 can use chemistry tools to access a variety of datasets, scientific papers and unstructured texts to extract and reason about material science information. The authors implemented a CoVe364 (Consistency Verification) scheme to evaluate the model’s answer and minimize hallucination. The authors showed that including CoVe increased the model’s precision by $\sim$ 20% when compared to previous methods such as an agent using ReAct only.369
Promoting molecular discovery is a topic with great attention in the literature devoted to it and, as described extensively above, LLMs have leveraged a large amount of unstructured data to accelerate that discovery. Janakarajan et al. 436 discuss the advantages of using LLMs in fields such as de novo drug design, reaction chemistry, and property prediction, but they augment the LLM in IBM ChemChat, a chatbot with the capability of using common APIs and python packages commonly used daily by a cheminformatics researcher to access molecular information. ChemChat has access to tools such as Generative Toolkit for Scientific Discovery (GT4SD)472, a package with dozens of trained models generative models for science, rxn4chemistry473, a package for computing chemistry reactions tasks, HuggingMolecules474, a package developed to aggregate molecular property prediction LMs, and RDKit475, a package to manipulating molecules. Since ChemChat implements an agent in a chat-like environment, users can interactively refine design ideas. Despite being developed to target de novo drug design, ChemChat nonetheless is a multi-purpose platform that can be more broadly used for molecular discovery.
In addition to the capabilities described above, LLM-based agents can empower users to tackle tasks that typically require extensive technical knowledge. In previous work, Wellawatte and Schwaller 286 and Gandhi and White 240 showed that including natural language explanations (NLE) in explainable AI (XAI) analysis can improve user understanding. More recently, Wellawatte and Schwaller 286 developed XpertAI286 to seamlessly integrate XAI techniques with LLMs to interpret and explain raw chemical data autonomously. Applying XAI techniques is usually restricted to technical experts but by integrating such techniques with an LLM-based agent to automate the workflow, the authors made XAI accessible to a wider audience.
Their system receives raw data with labels for physicochemical properties. The raw data is used to compute human-interpretable descriptors and then calculate SHAP values or Z-scores for Local Interpretable Model-agnostic Explanations (LIME). By calculating SHAP values (or SHapley Addictive exPlanations) a value can be assigned to each feature, indicating its contribution to a model’s output. LIME interprets a model by making a local approximation, around a particular prediction, to indicate what factors have contributed to that prediction in the model. It may use, for example, a surrogate local linear regression fit to recognized features.240 In addition to XAI tools, XpertAI can search and leverage scientific literature to provide accessible NLEs. While ChatGPT provides scientific justifications with similar accuracy, its explanation is often too broad. On the other hand, XpertAI provides data-specific explanations and visual XAI plots to support its explanations.286
These developments signify a growing trend in the integration of tools and LLMs in autonomous AI within scientific research. By automating routine tasks, enhancing information retrieval and analysis, and facilitating experimentation, AI is expanding the capabilities of researchers and accelerating the pace of scientific discovery. This review underscores the transformative impact of AI across various scientific domains, heralding a new era of innovation and efficiency in chemical research.
6 Challenges and opportunities
LLMs hold great potential in chemistry due to their ability to both predict properties and orchestrate existing computational and experimental tools. These capabilities enhance the accuracy and efficiency of chemical research and open up new avenues for discovery and innovation. By encapsulating AI models, data analysis software, and laboratory equipment within agent-based frameworks, researchers can harness these sophisticated tools through a unified interface. This approach not only simplifies the interaction with complex systems but also democratizes the immense capabilities of modern computational tools, thereby maximizing their utility in advancing chemical research and development.
Data Quality and Availability
Quality and availability of data are critical factors that influence the efficacy of LLMs. Indeed, scaling both the model size and the amount of training data used has proven to improve capabilities.476 However, current AI models are not trained on large amounts of chemical data, which limits their capabilities to reason about advanced chemical concepts.137
There are two types of datasets commonly used to train LLMs: unlabeled and labeled datasets. Unlabeled datasets, or pretraining data, are used during the semi-supervised training, which focuses on creating a “prior belief” about a molecule. Currently, we have huge datasets composed of hypothetical and/or theoretical data. When a model is trained on data that is not grounded in real chemical information, this might cause the model to learn a wrong prior belief.477
Labeled datasets, often used in benchmarks, also suffer from their inclusion of hypothetical and calculated data. Benchmarks are necessary for quantifying improvements in AI modeling and prediction within a competitive field. However, dominant benchmarks like MoleculeNet,55 have significant limitations that may restrict the generalizability and applicability of evolving models. In his blog, Walters 142 brings to light numerous errors and inconsistencies within the MoleculeNet data, which substantially impact model performance and reliability.138, 139, 140, 141. Walters also argues that the properties present in these benchmarks do not directly correlate with real chemistry improvement. As such, new benchmarks need to translate to practical chemistry problems directly. For instance, increasing accuracy in predicting LogP is not necessarily mapped to drugs with greater bioavailability. Some promising work has come from the Therapeutic Common Data (TDC)145, 144 includes data from actual therapeutic essays, providing a more practical foundation for model training.
The community continues to work to organize and curate datasets to make data ready for LLMs training and evaluation. Scientific benchmarks163, 430, 478, repositories with curated datasets136, and packages for model evaluation137 have been developed. However, the challenges concerning grounded truth and consistent datasets remain. With advancements in scientific document processing,316 there is now the opportunity to obtain new datasets from peer-reviewed scientific papers. Due to the multi-modal capabilities of such AI models, these new benchmarks can comprise multiple data types, potentially enhancing the applicability and transferability of these models. The continual curation of new, relevant datasets that represent the complexities of real-world chemical problems will further enhance the robustness and relevance of LLMs in chemistry.
Model Interpretability
Model interpretability is a significant challenge for LLMs due to their “black-box” nature, which obscures the understanding of how predictions are made. However, innovative approaches are being developed to enhance LLMs’ interpretability. For instance, Schwaller et al. 479 and Schilter et al. 480 used information from the different multi-attention heads. While Schwaller et al. 479 connected atoms from reactants to atoms in the products, Schilter et al. 480 assigned H-NMR peaks to specific hydrogens in a molecule to indicate how spectra were comprehended, or structures deduced. Additionally, since the LLMs use language, which is intrinsically interpretable, LLMs may be incrementally modified to explain their reasoning processes directly, exemplified with tools like eXpertAI286 and or simply adjusting prompting367, 415. These methods address the critical need for transparency in the mechanism of understanding for a good prediction beyond the good prediction itself.
Integration with Domain Knowledge and Cross-disciplinary Applications
While LLMs excel at pattern recognition, integrating explicit chemical rules and domain knowledge into these systems remains challenging. This integration is essential to make predictions that are not only statistically valid but also chemically reasonable. It was shown by Beltagy et al. 162 and Gu et al. 161 that better performance on common NLP tasks can be achieved by developing a vocabulary and pretraining on a domain-specific training corpus. While, pretrainng with domain-specific datasets that include chemical properties, reaction mechanisms, and experimental results may better capture the nuances of chemistry, the challenge to use AI to promote multi-disciplinary research remains. The Galactica LLM 155 also used special tokens for delineating chemical information, to relatively good success on chemistry tasks. Aryal et al. 481 also progress by creating an ensemble of specialist agents with different domains of knowledge, allowing them to interact to better answer the user query. Specifically, Aryal et al. 481 used agents with chemistry, physics, electrochemistry, and materials knowledge.
Tool development
The effectiveness of a combined LLM/autonomous agent approach hinges significantly on the availability and quality of the tools, as well as on the complexity and diversity of the chemical tasks at hand. Some emphasis should be placed on refining standalone tools, with the confidence that overarching frameworks, like a GPT-4-type wrapper, or “assistant”, will eventually integrate these tools seamlessly. Developers should stay informed about existing tools and design their tools to interface effectively with such a wrapper. This ensures that each tool is ready to contribute its unique capabilities to a cohesive agent system.
Reinforcement learning
RL has been successfully used in LLMs482, 483, 87, with a few applications also proposed for use in agents484, 485. The next frontier is applying RL to agents directly, to improve their ability on specific tasks. Bou et al. 96 provided a recent framework and example for generative molecular design when viewed as an RL problem (similar to RLHF) and some early success has been seen in applying the RLHF algorithm directly to protein language models where the reward model comes from scientific tasks.97 Neither of these are direct RL on language model agents, but are a step towards this goal.
Agent evaluation
Comparing different agent systems is challenging due to the lack of robust benchmarks and evaluation schemes. Consequently, it is difficult to define what constitutes a “superhuman” digital chemist and reach a consensus on the criteria for success. This issue is similar to the ongoing discussions about defining artificial general intelligence (AGI) and the expected capabilities of cognitive architectures.486, 487 Once a reliable metric for evaluating such AI systems is established, it is crucial for the AI scientific community to set clear guidelines for conducting research. Currently, assessing success is challenging because the goals are not well defined. Building on this, we propose using Bloom’s taxonomy as a reference point for developing a metric to evaluate more complex reasoning and tool use in autonomous agents. This educational framework categorizes cognitive skills in a hierarchical manner, from basic recall to creative construction, providing a structured approach to assess higher-order thinking and reasoning capabilities in these systems. This adaptation could significantly enhance the evaluation of LLMs and autonomous agents, especially when tackling complex chemical challenges.
Ethical and Safety Concerns
As with all AI technologies, there are ethical considerations in deploying LLMs, including bias in model predictions and the potential misuse of AI-generated chemical knowledge. Ruan et al. 488 and Tang et al. 461 pointed out the necessity of different levels of regulation in agent development, noting that current alignment methods may be insufficient for ensuring safety, and using human evaluation to address the risks of such systems is not scalable. Therefore, the lack of specialized models for risk control and effective safety evaluations poses perhaps the greatest challenges when assuring the safety of LLMs with the ability to use tools. This underscores the urgency, and the opportunity, in automating red-teaming of agents to enhance safety protocols in AI technologies.
Human-AI Collaboration in Chemical Research
LLMs are poised to transform fields such as drug discovery, materials science, and environmental chemistry due to their ability to predict chemical properties and reactions with remarkable accuracy. Models based on architectures like BERT have demonstrated their capability to achieve state-of-the-art performance in various property prediction tasks.43, 182 Furthermore, studies by Jablonka et al. 275 and Born and Manica 489 have showcased the predictive power of LLMs by reformulating traditional regression and classification tasks as generative tasks, opening up new avenues for chemical modeling. However, as emphasized by Weng 351, maintaining the reliability of LLM outputs is essential, as inaccuracies in formatting, logical reasoning, or content can significantly impede their practical utility. Hallucination is also an intrinsic issue with LLMs.490 Though agents can deal with hallucinations to some extent by implementing sanity-checking tools, it does not make the response hallucination-proof. A possible approach to address this issue is to use a human-in-the-loop approach, where steps of human-agent interaction are added to the workflow to check if the agent is in the correct pathway to solve the request.491, 492, 493
The potential of LLMs to design novel molecules and materials was highlighted by the AI-powered robotic lab assistant, A-Lab, which synthesized 41 new materials within just 17 days.494 Nonetheless, this achievement has sparked debates about the experimental methods and the actual integration of atoms into new crystalline materials, raising questions about the authenticity of the synthesized structures.495 These controversies underline the necessity for rigorous standards and the critical role of human expertise in validating AI-generated results. Again, the integration of advanced AI tools with the oversight of seasoned chemists is crucial, suggesting that a hybrid approach could significantly enhance both the innovation and integrity of materials science research.
In parallel, we have seen how LLM-based agents are increasingly capable of automating routine tasks in chemical research, which traditionally consume significant time and resources. These models excel in real-time data processing, managing vast datasets, and even conducting comprehensive literature reviews with minimal human intervention. Advances in AI technology now allow agents not only to perform predefined tasks but also to adapt and develop new tools for automating additional processes. For instance, tasks such as data analysis, literature review, and elements of experimental design are now being automated.286, 496, 497, 430, 435, 433, 45, 46 This automation liberates chemists to focus on more innovative and intellectually engaging aspects of their work, and the opportunity is to expand productivity and creativity in their science.
7 Conclusions
Since this review is targeted in part to an audience of chemists, who may not have yet embraced AI technology, we consider it valuable to point out our perspective that AI in chemistry is definitely here to stay. We predict that its use will only grow as a necessary tool that will inevitably lead to more jobs and greater progress. We hope to facilitate the change by connecting the technology to the chemical problems that our readership is already addressing through more traditional methods.
Large Language Models (LLMs) have demonstrated remarkable potential in reshaping chemical research and development workflows. These models have facilitated significant advancements in molecular simulation, reaction prediction, and materials discovery. In this review, we discussed the evolution of LLMs in chemistry and biochemistry. Successful cases where LLMs have proven their potential in promoting scientific discovery were shown with caveats of such models.
Adopting LLM-based autonomous agents in chemistry has enhanced the accuracy and efficiency of traditional research methodologies and introduced innovative approaches to solving complex chemical problems. Looking forward, the continued integration of LLMs promises to accelerate the field’s evolution further, driving forward the frontiers of scientific discovery and technological innovation in chemistry. We have shown how agents have been used in chemistry and proposed a framework for thinking about agents as a central LLM followed by interchangeable components.
However, despite the community’s astonishing advances in this field, many challenges still require solutions. We identified the main challenges and opportunities that need to be addressed to promote the further development of agents in chemistry. Addressing the challenges related to model transparency, data biases, and computational demands will be crucial for maximizing their utility and ensuring their responsible use in future scientific endeavors.
While there are significant challenges to be addressed, the opportunities presented by LLMs in chemistry are vast and have the potential to fundamentally alter how chemical research and development are conducted. Effectively addressing these challenges will be crucial for realizing the full potential of LLMs in this exciting field. To keep pace with the ever-growing, we will maintain a repository with an organized structure listing new studies regarding LLMs and LLM-based agents focused on scientific purposes. The repository can be found in https://github.com/ur-whitelab/LLMs-in-science
Author contribution
All authors contributed to writing this review article.
Competing interests
The authors have no conflicts to declare.
Acknowledgments
M.C.R. and A.D.W. acknowledge the U.S. Department of Energy, Grant No. DE-SC0023354, and C.J.C. gratefully acknowledges the Jane King Harris Endowed Professorship at Rochester Institute of Technology for the support provided for this publication. We are grateful for feedback from early drafts of this review from the following colleagues: Kevin Jablonka, Philippe Schwaller, Michael Pieler, and Ryan-Rhys Griffiths.